### v3

# Composite Learning Backstepping Control With Guaranteed Exponential Stability and Robustness

Abstract
Adaptive backstepping control provides a feasible solution to achieve asymptotic tracking for mismatched uncertain nonlinear systems. However, the closed-loop stability depends on high-gain feedback generated by nonlinear damping terms, and closed-loop exponential stability with parameter convergence involves a stringent condition named persistent excitation (PE). This paper proposes a composite learning backstepping control (CLBC) strategy based on modular backstepping and high-order tuners to compensate for the transient process of parameter estimation and achieve closed-loop exponential stability without the nonlinear damping terms and the PE condition. A novel composite learning mechanism is designed to maximize the staged exciting strength for parameter estimation, such that parameter convergence can be achieved under a condition of interval excitation (IE) or even partial IE that is strictly weaker than PE. An extra prediction error is employed in the adaptive law to ensure the transient performance without nonlinear damping terms. The exponential stability of the closed-loop system is proved rigorously under the partial IE or IE condition. Simulations have demonstrated the effectiveness and superiority of the proposed method in both parameter estimation and control compared to state-of-the-art methods.
I Introduction
Adaptive control is desirable due to its unique capacity to accommodate uncertain and time-varying properties of nonlinear systems, where recent survey papers can be referred to [31, 6, 42, 3, 46, 41]. The presence of mismatched uncertainties is a major obstacle to adaptive control of nonlinear systems. Adaptive integral backstepping with overparameterization, which combines integral backstepping and direct adaptive control, is a precursor to relax the above obstacle by designing an adaptive law to adjust a virtual control input at each backstepping step [18]. A tuning function approach is a direct adaptive backstepping approach without overparameterization [84, Ch. 4], where an adaptive law termed as a tuning function is constructed iteratively at each backstepping step, while an actual adaptive law is generated at the last step by all previous tuning functions. It is revealed that adaptive backstepping control driven by tuning functions has a higher-order tracking property [32]. There exist two common drawbacks for the above adaptive backstepping approaches: 1) The “explosion of complexity” exists due to the repeated differentiation of virtual control inputs; 2) the exponential stability of the closed-loop system (implying parameter convergence and robustness) relies on a strict condition termed persistent excitation (PE), which requires system states contain sufficiently rich spectral information all the time.
Adaptive dynamic surface control (DSC) applies a first-order linear filter to estimate the time derivative of each virtual control input at its backstepping step for addressing the complexity problem of the classical backstepping control [49]. The performance and robustness of adaptive DSC have been enhanced by integrating neural network (NN) approximation [70], prescribed performance control [4], nonlinear filtering [79], coordinate transformation [77], power integration [26], etc. An approach similar to DSC, named command-filtered backstepping control (CFBC), employs second-order linear filters to estimate the time derivatives of virtual control inputs and compensation terms for stability guarantees [14]. The performance and robustness of adaptive CFBC have been improved by incorporating NN approximation [15], exact differentiation [47], immersion and invariance [16], etc. Yat, adaptive DSC and CFBC suffer from the explosion of dynamic order and the loss of global stability and asymptotic tracking due to filtering operations [50].
A modular backstepping approach follows the certainty equivalence principle that separates control and estimation designs [84, Ch. 5]. A key feature of this approach is that the time derivatives of virtual control inputs are replaced by their partial derivatives with respect to system states and reference signals, while the resulting high-order time derivatives of parameter estimates are treated as additive disturbances. Thus, the modular backstepping approach does not involve tuning functions or overparameterization and has lower complexity. This approach ensures closed-loop stability with strong robustness by introducing a nonlinear damping term in a stabilizing function at each backstepping step [71]. A standard gradient-descent identifier derived from a swapping scheme can be combined with modular backstepping to achieve asymptotic tracking [84, Ch. 6]. Nevertheless, as the high-order time derivatives of parameter estimates exist in the closed-loop system, the modular backstepping approach may degrade transient tracking and prevent exact parameter estimation even in the presence of the PE condition.
A high-order tuner (HOT) approach can efficiently remove the negative influence caused by the time derivatives of parameter estimates in modular backstepping [45]. The key idea of the HOT is to apply a linear filter with a sufficiently high relative degree to the adaptive law, such that the exact implementation of the high-order time derivatives of parameter estimates becomes feasible. In [30], a direct adaptive control scheme was combined with the HOT to counteract the transient process of parameter estimates caused by their high-order time derivatives, improving the transient and steady-state tracking performance. In [43], a memory regressor extension (MRE) identifier, which utilizes regressor extension with filtering, was combined with the HOT to design an indirect adaptive control law, where the HOT is applied to an extended regression model such that the high-order time derivatives of parameter estimates can be calculated exactly without filtering delay. However, the above two methods still need nonlinear damping terms to ensure closed-loop stability and transient performance and resort to the stringent PE condition for exponential stability guarantees.
From the above discussions, existing modular backstepping methods have the following limitations:
  1. 1.
    The transient and steady-state tracking performances rely on nonlinear damping terms;
  2. 2.
    The transient process of parameter estimates due to their high-order time derivatives can destroy the tracking performance and parameter convergence;
  3. 3.
    The stringent PE condition must be fulfilled to realize the exponential stability of the closed-loop system.
Motivated by the above facts, this paper proposes a composite learning backstepping control (CLBC) strategy that ensures the exponential stability of the closed-loop system under relaxed excitation conditions for strict-feedback uncertain nonlinear systems. The design procedure is as follows: First, the modular backstepping scheme without nonlinear damping is introduced to facilitate the control design; second, a generalized regression equation is constructed by the swapping technique with interval integrations; third, a linear filter is applied to the generalized regression equation to generate a linearly parameterized model; fourth, a generalized prediction error is designed to exploit online data memory; fifth, a general prediction error is introduced to counteract a modeling error term; finally, a composite learning HOT is constructed by combining the two prediction errors to implement the high-order time derivatives of parameter estimates exactly. The contributions of this study lie in threefold:
  1. 1.
    A feasible modular backstepping strategy termed CLBC is proposed to guarantee transient and steady-state tracking without nonlinear damping terms or high control gains;
  2. 2.
    An algorithm of staged exciting strength maximization is designed to enhance the online data memory of composite learning in different partial excitation stages;
  3. 3.
    The exponential stability and robustness of the closed-loop system with parameter convergence are proven under the condition of interval excitation (IE) or even partial IE.
Notations: $\min\{\cdot\}$ denote the minimum operator, $\lambda_{\min}(A)$ and $\lambda_{\max}(A)$ are the minimal and maximum eigenvalues of $A$, respectively, $\sigma_{\min}(A)$ is the minimum singular value of $A$, $\|\bm{x}\|$ is the Euclidean norm of $\bm{x}$, $L_{\infty}$ is the space of bounded signals, $I$ is an identity matrix, $\bm{0}$ is a zero matrix, $\Omega_{c}$ :=:= $\{\bm{x}|\|\bm{x}\|\leq c\}$ is the ball of radius $c$, $\arg\max_{x\in S}f(x)$ :=:= $\{x\in S|f(y)$ $\leq$ $f(x)$, $\forall y$ $\in$ S}S\}, $\bm{g}\in\mathcal{C}^{k}$ indicates that $\bm{g}$ has continuous partial derivatives up to the order $k$, where $A\in\mathbb{R}^{n\times n}$, $\bm{x}\in\mathbb{R}^{n}$, $c\in\mathbb{R}^{+}$, $f:\mathbb{R}\mapsto\mathbb{R}$, $\bm{g}$ :: $\mathbb{R}^{n}\mapsto\mathbb{R}^{m},$ $S\subset\mathbb{R}$, $n,m$ are positive integers, and $k$ is an non-negative integer.
II Problem Formulation
Consider a class of $n$th-order strict-feedback uncertain nonlinear systems as follows [84] [^1] This study considers the system with linear-in-the-parameters uncertainties in (4), but the following theoretical results can be extended to certain systems with nonlinear-in-the-parameters uncertainties as discussed in [59].:
$\left\{\begin{array}[]{l}\dot{x}_{i}={\bm{\varphi}}_{i}^{T}({\bm{x}}_{i})\bm{\theta}+x_{i+1},\\ \dot{x}_{n}={\bm{\varphi}}_{n}^{T}(\bm{x})\bm{\theta}+\beta(\bm{x})u,\\ y=x_{1}\end{array}\right$. (4)
with $i=$ 1 to n1n-1 and ${\bm{x}}_{i}(t):=[x_{1}(t),x_{2}(t),\cdots,x_{i}(t)]^{T}\in\mathbb{R}^{i}$, where $\bm{x}(t):=[x_{1}(t),x_{2}(t),\cdots,x_{n}(t)]^{T}$ $\in\mathbb{R}^{n}$ is a system state, $u(t)\in\mathbb{R}^{n}$ is a control input, $y(t)\in\mathbb{R}$ is a system output, $\bm{\theta}\in\Omega_{{\rm c}_{\theta}}\subset\mathbb{R}^{N}$ with ${\rm c}_{\theta}\in\mathbb{R}^{+}$ is a unknown parameter vector, ${\bm{\varphi}}_{i}:\mathbb{R}^{i}$ $\rightarrow\mathbb{R}^{N}$ denotes a known regressor, $\beta:$ $\mathbb{R}^{n}\rightarrow\mathbb{R}$ is a known gain function, and $N$ is the number of parameter elements. Let $y_{\rm r}(t)\in\mathbb{R}$ be a reference signal. The following definitions are given for the subsequent analysis.
Definition 1 [83]: A bounded regressor $\Phi(t)\in\mathbb{R}^{N\times n}$ is of PE if there exist constants $t_{0}$, $\sigma$, and $\tau_{\rm d}\in\mathbb{R}^{+}$ such that
$\int_{t-\tau_{\rm d}}^{t}\Phi(\tau)\Phi^{T}(\tau)d\tau\geq\sigma I,\forall t\geq t_{0}.$
Definition 2 [19]: A bounded regressor $\Phi(t)\in\mathbb{R}^{N\times n}$ is of IE if there exist constants $\sigma$, $\tau_{\rm d}$, and $T_{\rm e}\in\mathbb{R}^{+}$ such that
$\int_{T_{\rm e}-\tau_{\rm d}}^{T_{\rm e}}\Phi(\tau)\Phi^{T}(\tau)d\tau\geq\sigma I.$
Definition 3: A bounded regressor $\Phi(t)\in\mathbb{R}^{N\times n}$ is of partial IE, if there exist constants $\sigma$, $\tau_{\rm d}$, and $T_{\rm a}$ $\in$ $\mathbb{R}^{+}$ such that
$\int_{T_{\rm a}-\tau_{\rm d}}^{T_{\rm a}}\Phi_{\zeta}(\tau)\Phi_{\zeta}^{T}(\tau)d\tau\geq\sigma I$
in which $\Phi_{\zeta}\in\mathbb{R}^{m\times n}$ is a sub-regressor obtained by eliminating some rows of $\Phi$ with $1\leq m<N$.
For convenience, a column ${\bm{\phi}}_{j}(t)\in\mathbb{R}^{n}$ ($j=1$ to $N$) of a regressor $\Phi^{T}(t)\in\mathbb{R}^{n\times N}$ is named as a channel. Consequently, one has $\Phi(t)$ == ${\bm{\phi}}_{1}(t),{\bm{\phi}}_{2}(t),\cdots,{\bm{\phi}}_{N}(t)]^{T}$. A channel ${\bm{\phi}}_{j}(t)$ is named as an active channel if $\|{\bm{\phi}}_{j}(t)\|\neq 0$, conversely termed as an inactive channel. Without loss of generality, assume that there exists a proper time window $\tau_{\rm d}$ satisfying either Definition 2 or Definition 3, and consider the case where IE may not exist, $\forall t\geq 0$, but partial IE exists at the beginning and some moments later. We aim to design a suitable adaptive control strategy for the system (4) such that closed-loop stability with parameter convergence can be guaranteed without the PE condition.
Remark 1: The PE and IE conditions require that all channels $\bm{\phi}_{j}(t)$ ($j=1$ to $N$) are activated in an uncorrelated manner within a time window $t-\tau_{\rm d},t]$ (sliding in the case of PE), which is difficult to satisfy in many practical scenarios due to the presence of inactive channels (i.e., there exists $j\in\{1,2,\cdots,N\}$ such that $\|\bm{\phi}_{j}(\tau)\|=0$, $\forall\tau\in[t-\tau_{\rm d},t]$). The partial IE condition relaxes the requirement by ignoring all inactive channels during $t-\tau_{\rm d},t]$, which allows it to satisfy at the beginning and some moments later due to the changing state $\bm{x}(t)$ over time in general cases. This implies the inexistence of the case with all channels $\|\bm{\phi}_{j}(t)\|\equiv 0$ ($j=1$ to $N$), $\forall t\geq 0$.
III Modular Backstepping Control Design
The following assumptions presented in [84] are given for the modular backstepping control design.
Assumption 1: $y_{\rm r}$, $\dot{y}_{\rm r}$, $\cdots$, $y^{(n-1)}_{\rm r} \in$ $L_{\infty}$, $\beta(\bm{x})\neq 0$, $\forall\bm{x}\in\mathbb{R}^{n}$, and $\bm{\varphi}_{i}\in\mathcal{C}^{n-i}$ for $i=1$ to $n$.
The virtual control inputs $v_{1}(t),v_{i}(t)\in\mathbb{R}$ in the modular backstepping approach are recursively given by [84]
$v_{1}(x_{1},\hat{\bm{\theta}},y_{\rm r})=-k_{{\rm c}1}e_{1}-\bm{\psi}_{1}^{T}\hat{\bm{\theta}}$, (5a)
$v_{i}({\bm{x}}_{i},\Theta_{i-1},{\bm{y}}_{{\rm r}i})=-k_{{\rm c}i}e_{i}-e_{i-1}-\bm{\psi}_{i}^{T}\hat{\bm{\theta}}$
$+$ $\sum_{k=1}^{i-1}\left(\frac{\partial v_{i-1}}{\partial x_{k}}x_{k+1}+\frac{\partial v_{i-1}}{\partial\hat{\bm{\theta}}^{(k-1)}}\hat{\bm{\theta}}^{(k)}+\frac{\partial v_{i-1}}{\partial y_{\rm r}^{(k-1)}}y_{\rm r}^{(k)}\right)$ (5b)
with $\Theta_{i-1}(t)$ :=:= $\big[\hat{\bm{\theta}}(t),\dot{\hat{\bm{\theta}}}(t),\cdots,\hat{\bm{\theta}}^{(i-1)}(t)\big]$ $\in\mathbb{R}^{N\times i}$ and ${\bm{y}}_{{\rm r}i}(t)$ :=:= $\big[y_{\rm r}(t)$, $\dot{y}_{\rm r}(t)$, $\cdots$, $y^{(i-1)}_{\rm r}(t)\big]^{T}$ $\in\mathbb{R}^{i}$, where $e_{1}(t)$ :=:= $x_{1}(t)-y_{\rm r}(t)\in\mathbb{R}$ and $e_{i}(t)$ :=:= $x_{i}(t)-v_{i-1}(t)-y_{\rm r}^{(i-1)}(t)\in\mathbb{R}$ are tracking errors, $\bm{\psi}_{1}:=\bm{\varphi}_{1}$ and $\bm{\psi}_{i}:=\bm{\varphi}_{i}-\sum_{k=1}^{i-1}\frac{\partial v_{i-1}}{\partial x_{k}}\bm{\varphi}_{k}\in\mathbb{R}^{N}$ are regressors, $k_{{\rm c}1},k_{{\rm c}i}\in\mathbb{R}^{+}$ are control gain parameters, and $i=2,\cdots,n$. In the final step, design the control law
$u=\frac{1}{\beta(\bm{x})}\left(v_{n}(\bm{x},\Theta_{n-1},{\bm{y}}_{{\rm r}n})+y_{\rm r}^{(n)}\right).$ (6)
Applying (5), (6), and $e_{i}$ == $x_{i}-v_{i-1}-y_{\rm r}^{(i-1)}$ to (4) results in a closed-loop tracking error system
$\dot{\bm{e}}=\Lambda\bm{e}+\Phi^{T}(\bm{x},\Theta_{n-1},{\bm{y}}_{{\rm r}n})\tilde{\bm{\theta}}$ (7)
in which $\bm{e}(t)$ :=:= $e_{1}(t),e_{2}(t),\cdots,e_{n}(t)]^{T}\in\mathbb{R}^{n}$ is a tracking error vector, $\Phi$ :=:= $\bm{\psi}_{1}$, $\bm{\psi}_{2}$, $\cdots$, $\bm{\psi}_{n}]$ $\in\mathbb{R}^{N\times n}$ is a new regressor, $\tilde{\bm{\theta}}(t)$ :=:= $\bm{\theta}-\hat{\bm{\theta}}(t)$ $\in\mathbb{R}^{N}$ is a parameter estimation error, and
$\Lambda=\left[\begin{matrix}-k_{{\rm c}1}&1&0&\cdots&0\\ -1&-k_{{\rm c}2}&1&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&\cdots&1\\ 0&0&0&\cdots&-k_{{\rm c}n}\\ \end{matrix}\right]\in\mathbb{R}^{n\times n}.$
Consider linear filtering operations
$\left\{\begin{array}[]{ll}\dot{\bm{\zeta}}(t)=\Lambda\bm{\zeta}(t)+\Phi^{T}(t)\hat{\bm{\theta}}(t),&\bm{\zeta}(0)=-\bm{e}(0)\\ \dot{\Phi}_{\rm s}^{T}(t)=\Lambda{\Phi}_{\rm s}^{T}(t)+\Phi^{T}(t),&{\Phi}_{\rm s}(0)=\bm{0}\end{array}\right$. (10)
and define an output vector $\bm{p}(t)$ :=:= $\bm{e}(t)+\bm{\zeta}(t)\in\mathbb{R}^{n}$, where $\bm{\zeta}(t)\in\mathbb{R}^{n}$ and ${\Phi}_{\rm s}(t)\in\mathbb{R}^{N\times n}$ are filtered outputs. Following the swapping technique [84, Ch. 6] and (10), one can obtain a static linear parametric model as a regression equation as follows:
$\bm{p}(t)=\Phi^{T}_{\rm s}(t)\bm{\theta}.$ (11)
However, the parameter vector $\bm{\theta}$ in (11) can not be estimated by classical adaptation schemes because the regressor $\Phi_{\rm s}$ relies on the inaccessible high-order time derivatives $\hat{\bm{\theta}}^{(k)}$ ($k=1$ to n1n-1). Even when $\hat{\bm{\theta}}^{(k)}$ are available, the parameter convergence of classical adaptation schemes relies on the stringent PE condition, which requires that the reference trajectory $y_{\rm r}$ includes sufficiently rich spectral information all the time.
Remark 2: In existing modular backstepping control methods, nonlinear damping terms $k_{\rm{d}\it i}\|\bm{\psi}_{i}\|^{2}$ must be incorporated into the virtual control inputs $v_{i}$ in (5) to counteract the transient response arising from the modeling error term $\Phi^{T}\tilde{\bm{\theta}}$ and the high-order time derivatives $\hat{\bm{\theta}}^{(k)}$ ($k=1$ to n1n-1), so the boundedness of the closed-loop system can always be established even in the absence of adaptation, where $k_{\rm{d}\it i}\in\mathbb{R}^{+}$ are damping parameters ($i=1$ to $n$) [84, 43]. However, the square terms $\|\bm{\psi}_{i}\|^{2}$ can lead to high-gain control, regardless of whether the estimation error $\tilde{\bm{\theta}}$ converges to $\bm{0}$ or not, which results in noise amplification and control saturation in practice. In this study, we aim to establish closed-loop stability without resorting to nonlinear damping terms as detailed in Sec. IV.
IV Composite Learning Design
IV-A Composite Learning High-Order Tuner
For convenience, let $\Phi_{{\rm s},\zeta}$ :=:= $\bm{\phi}_{{\rm s},k_{1}},\bm{\phi}_{{\rm s},k_{2}},\cdots,\bm{\phi}_{{\rm s},k_{N_{\zeta}}}]^{T}$ $\in\mathbb{R}^{N_{\zeta}\times n}$ ($1\leq k_{j}\leq N$, $j=1$ to $N_{\zeta}$) be an active sub-regressor of $\Phi_{\rm s}$ in (11), $N_{\zeta}<N$ be the number of active channels, and $\bm{\psi}_{\zeta,i}$ $\in\mathbb{R}^{N_{\zeta}}$ ($i=1$ to $n$) be the $i$th column of $\Phi_{{\rm s},\zeta}$. The following partial identifiability assumption is introduced to facilitate the composite learning HOT design and to ensure the existence of partial IE at the beginning and some moments later.
[…]
Multiplying (11) by $\Phi_{\rm s}$ and letting $\bm{\zeta}(0)=-\bm{e}(0)$ and ${\Phi}_{\rm s}(0)=\bm{0}$, one obtains an extended regression equation
$\Phi_{\rm s}(t)\Phi_{\rm s}^{T}(t)\bm{\theta}=\Phi_{\rm s}(t)\bm{p}(t).$ (12)
Integrating (12) over a moving time window $t-\tau_{\rm d},t]$, one obtains a generalized regression equation
$\Psi(t)\bm{\theta}=\bm{q}(t)$ (13)
where $\Psi(t)\in\mathbb{R}^{N\times N}$ is an excitation matrix given by
$\Psi(t):=\int_{t-\tau_{\rm d}}^{t}\Phi_{\rm s}(\tau)\Phi_{\rm s}^{T}(\tau)d\tau$ (14)
and $\bm{q}(t)\in\mathbb{R}^{N}$ is an auxiliary variable given by
$\bm{q}(t):=\int_{t-\tau_{\rm d}}^{t}\Phi_{\rm s}(\tau)\bm{p}(\tau)d\tau.$ (15)
To obtain the high-order time derivatives $\hat{\bm{\theta}}^{(k)}$, it is feasible to apply a linear filter with n1n-1 relative degrees
$H(s):=\prod_{i=1}^{n-1}\frac{\alpha_{i}}{s+\alpha_{i}}$ (16)
with $\alpha_{i}\in\mathbb{R}^{+}$ ($i=$ 1 to $n$) being filtering constants to (13), which results in a generalized parameterized model
$Q(t)\bm{\theta}={\bm{q}}_{\rm f}(t)$ (17)
with $Q(t):=H(s)[\Psi(t)]$ and ${\bm{q}}_{\rm f}(t):=H(s)[\bm{q}(t)]$. Then, an adaptive law of $\hat{\bm{\theta}}$ can be designed such that $\hat{\bm{\theta}}^{(k)}$ is obtainable by the direct differentiation of filtered elements on $\Psi$ and $\bm{q}$.
From Assumption 2, partial IE exists at the beginning and some moments later, and there exist $\sigma$, $\tau_{\rm d}\in\mathbb{R}^{+}$ to get
$\Psi_{\zeta}(t):=\int_{t-\tau_{\rm d}}^{t}\Phi_{\rm s,\zeta}(\tau)\Phi_{\rm s,\zeta}^{T}(\tau)d\tau\geq\sigma I$ (18)
in which $\Phi_{\rm s,\zeta}\in\mathbb{R}^{N_{\zeta}\times n}$ is a sub-regressor composed of all active channels ${\bm{\phi}}_{{\rm s},k_{j}}$ of $\Phi_{\rm s}$, i.e., $\Phi_{\rm s,\zeta}:=[{\bm{\phi}}_{{\rm s},k_{1}},{\bm{\phi}}_{{\rm s},k_{2}},\cdots,{\bm{\phi}}_{{\rm s},k_{N_{\zeta}}}]^{T}$ with $\|{\bm{\phi}}_{{\rm s},k_{j}}(\tau_{j})\|>0$, $\exists\tau_{j}\in[t-\tau_{\rm d},t]$, $1\leq k_{j}\leq N$, and $j=1$ to $N_{\zeta}$. If the IE condition holds, then there exists a finite time $T_{\rm e}\in\mathbb{R}^{+}$ such that $\Psi(T_{\rm e})\geq\sigma I$; otherwise, $T_{\rm e}=\infty$.
The index $k_{j}$ of active channels $\bm{\phi}_{{\rm s},k_{j}}$ may be changed under partial IE, which leads to the existence of multiple partial IE stages. To consider the changes of the sub-regressor $\Phi_{{\rm s},\zeta}$ under different partial IE stages, let $\mathcal{I}$ :=:= $\{k_{1},k_{2},\cdots,k_{N_{\zeta}}\}$ and $\mathcal{I}^{\prime}$ :=:= $\{k_{1}^{\prime},k_{2}^{\prime},\cdots,k_{N_{\zeta}^{\prime}}^{\prime}\}$ ($1\leq k_{j}^{\prime}\leq N$ and $j=1$ to $N_{\zeta}^{\prime}$) be index sets of active channels in the current and previous partial IE stages, respectively, where $N_{\zeta}^{\prime}<N$ is the number of previous active channels. Then, Algorithm 1 is provided to reconstruct the sub-regressor $\Phi_{{\rm s},\zeta}$ and maximize the exciting strength $\sigma_{\min}(\Psi_{\zeta}(t))$ in each partial IE stage, where $T_{\rm s}\in\mathbb{R}^{+}$ is a sampling time, $T_{\rm a}\in\mathbb{R}^{+}$ is the first epoch in each partial IE stage, $\Psi_{k_{j},k_{j}}(t)$ :=:= $\int_{t-\tau_{\rm d}}^{t}\|\bm{\phi}_{{\rm s},k_{j}}(\tau)\|^{2}d\tau$ is the $k_{j}$th diagonal element of $\Psi(t)$, $\sigma_{\rm c}(t)\in\mathbb{R}^{+}$ is the current maximal exciting strength, and $t_{\rm e}\in\mathbb{R}^{+}$ is the corresponding exciting time. Based on the above argument, define a generalized prediction error
$\bm{\xi}(t):={\bm{q}}_{\rm f}(t,t_{\rm e})-Q(t,t_{\rm e})\hat{\bm{\theta}}(t)$ (19)
with $Q(t,t_{\rm e})$ :=:= $H(s)[\Psi(t_{\rm e})]$ and ${\bm{q}}_{\rm f}(t,t_{\rm e})$ :=:= $H(s)[{\bm{q}}(t_{\rm e})]$.
Since the system (4) is continuous-time and Algorithm 1 is discrete-time implemented with the sampling time $T_{\rm s}$, $\tau_{\rm d}$ should be chosen to be greater than $T_{\rm s}$ to ensure the correct functioning of Algorithm 1. In Algorithm 1, we first choose a sufficiently small threshold $\sigma$ [see Line 1 in Algorithm 1]. Note that partial IE usually exists; otherwise, all channels will be deactivated. At the beginning of each partial IE stage, the maximal exciting strength $\sigma_{\rm c}$ is reset to $\sigma$ [see Line 7 in Algorithm 1]. If the current exciting strength $\sigma_{\min}(\Psi_{\zeta}(t))$ is greater than $\sigma_{\rm c}$, $t_{\rm e}$ and $\sigma_{\rm c}$ are updated [see Lines 9–11 in Algorithm 1]; otherwise, they remain unchanged. Thus, Algorithm 1 ensures that the exciting strength $\sigma_{\min}(\Psi_{\zeta}(t_{\rm e}))$ is monotonically non-decreasing at each partial IE stage. Fig. 1 illustrates $\sigma_{\rm c}$ in Algorithm 1 for a simple case with two partial IE stages. As new active channels $\bm{\phi}_{{\rm s},k_{j}}(t)$ exist at $t=T_{\rm a}$, the sub-regressor $\Phi_{{\rm s},\zeta}$ with the excitation matrix $\Psi_{\zeta}(t)$ in (18) is reconstructed by all new active channels [see Lines 4–8 in Algorithm 1]. In this partial IE stage, as the exciting strength $\sigma_{\min}(\Psi_{\zeta}(t))$ can be time-varying [see the green dash line in Fig. 1], $t_{\rm e}$ is updated based on exciting strength maximization, i.e., $\max_{\tau\in[T_{\rm a},t]}\sigma_{\min}(\Psi_{\zeta}(\tau))$ [see Lines 9–11 in Algorithm 1 and the black solid line in Fig. 1]. This is the same for the IE stage [see Lines 13–15 in Algorithm 1 and the red area in Fig. 1].
Algorithm 1 Staged exciting strength maximization
1:Initialize: $\mathcal{I}^{\prime}\leftarrow\emptyset$, $T_{\rm a}=0$, $\sigma_{\rm c}=\sigma$, $t_{\rm e}=0$
2:for $t=0$ with a step size of $T_{\rm s}$ do
3:  if length()<N($\mathcal{I}^{\prime}$)<N then
4:   Find the indexes $k_{j}$ satisfying $\Psi_{k_{j},k_{j}}(t)>0$, $k_{j}\in\{1,2,\cdots,N\}$ and set $\mathcal{I}\leftarrow\{k_{1},k_{2},\cdots,k_{N_{\zeta}}\}$
5:   if $\exists k_{j}\in\mathcal{I}$ such that $k_{j}\notin\mathcal{I}^{\prime}$ then
6:     Reconstruct $\Phi_{\rm s,\zeta}(t)$ and $\Psi_{\zeta}(t)$ by $\mathcal{I}$
7:     $\sigma_{\rm c}\leftarrow\sigma$, $T_{\rm a}\leftarrow t$, $\mathcal{I}^{\prime}\leftarrow\mathcal{I}$
8:   end if
9:   if $\sigma_{\min}(\Psi_{\zeta}(t))\geq\sigma_{\rm c}$ then
10:     $\sigma_{\rm c}\leftarrow\sigma_{\min}(\Psi_{\zeta}(t))$, $t_{\rm e}\leftarrow t$
11:   end if
12:  else
13:   if $\sigma_{\min}(\Psi(t))\geq\sigma_{\rm c}$ then
14:     $\sigma_{\rm c}\leftarrow\sigma_{\min}(\Psi(t))$, $t_{\rm e}\leftarrow t$
15:   end if
16:  end if
17:end for
To counteract the transient process of the modeling error term $\Phi^{T}\tilde{\bm{\theta}}$ in (7) such that closed-loop stability can be ensured without resorting to the nonlinear damping terms $k_{\rm{d}\it i}\|\bm{\psi}_{i}\|^{2}$ ($i=$ 1 to $n$), we introduce extra prediction error feedback. Applying $H(s)$ in (16) to (7) results in a filtered regression equation
$\bm{z}(t)=\Phi_{\rm f}^{T}(t)\bm{\theta}$ (20)
with $\bm{z}(t)$ :=:= $sH(s)[\bm{e}]+H(s)[\Phi^{T}\hat{\bm{\theta}}-\Lambda\bm{e}]$ and $\Phi_{\rm f}$ :=:= $H(s)[\Phi]$. Then, giving a filtered prediction model
$\hat{\bm{z}}(t)=\Phi_{\rm f}^{T}(t)\hat{\bm{\theta}}$ (21)
define a general filtered prediction error
$\bm{\epsilon}(t):=\bm{z}(t)-\Phi_{\rm f}^{T}(t)\hat{\bm{\theta}}(t)$ (22)
where $\hat{\bm{z}}(t)\in\mathbb{R}^{n}$ is a predicted value of $\bm{z}(t)$.
With the generalized prediction error $\bm{\xi}(t)$ in (19) and the prediction error $\bm{\epsilon}(t)$ in (22), design a composite learning HOT
$\dot{\hat{\bm{\theta}}}=\Gamma\left(\Phi_{\rm f}\bm{\epsilon}(t)+\kappa\bm{\xi}(t)\right)$ (23)
in which $\Gamma\in\mathbb{R}^{N\times N}$ is a positive-definite diagonal matrix of learning rates, and $\kappa\in\mathbb{R}^{+}$ is a weighting factor. Since $H(s)$ in (16) owns n1n-1 relative degrees, the time derivatives of $\hat{\bm{\theta}}$ in (23) up to the (n1n-1)th order can be implemented physically by a direct differentiation scheme [35]. More specifically, the high-order time derivatives of $\hat{\bm{\theta}}$ are calculated by
$\hat{\bm{\theta}}^{(k+1)}=\Gamma\left(\sum_{i=0}^{k}C_{k}^{i}s^{k-i}H(s)[\Phi]\bm{\epsilon}^{(i)}+\kappa\bm{\xi}^{(k)}\right)$ (24)
with $\hat{\bm{\theta}}^{(i)}(0)$ == $\bm{0}$ and intermediate time derivatives
$\bm{\epsilon}^{(i)}=\bm{z}^{(i)}-\sum_{j=0}^{i}C_{i}^{j}s^{i-j}H(s)[\Phi]\hat{\bm{\theta}}^{(j)},$
$\bm{\xi}^{(k)}={\bm{q}}_{\rm f}^{(k)}(t,t_{\rm e})-\sum_{i=0}^{k}C_{k}^{i}Q^{(k-i)}(t,t_{\rm e})\hat{\bm{\theta}}^{(i)}$
in which $C_{k}^{i}=k!/(i!(k-i)!)$ are binomial coefficients with $0\leq i\leq k$ and $1\leq k\leq n-2$.
Figure 1: An illustration of the current maximal exciting strength $\sigma_{\rm c}$ in Algorithm 1. Note that the black solid line denotes $\sigma_{\rm c}$, the blue and green dash lines are $\sigma_{\min}(\Psi_{\zeta})$ in two partial IE stages, and the red dotted line is $\sigma_{\min}(\Psi)$.
Figure 2: An illustration of the relationship between the channels $\bm{\phi}_{{\rm s},i}$ and the excitation matrix $\Psi$ with $N=7$. Note that the blue modules denote active channels, the white modules are inactive channels, and $\bm{p}$ is defined in (11).
IV-B Some Meaningful Discussions
In the proposed composite learning HOT (23), the generalized prediction error $\bm{\xi}(t)$ in (19) is updated by a newly designed algorithm termed staged exciting strength maximization, which enables online data memory to be established under different partial IE stages, whereas online data memory in the classical composite learning can only be achieved under IE as its exciting strength is obtained directly by calculating the minimum singular value of the excitation matrix $\Psi(t)$ [50]. Note that IE requires all regressor channels to be activated in an uncorrelated manner at a time window $T_{\rm e}-\tau_{\rm d},T_{\rm e}]$, but partial IE only requires partial regressor channels to be activated at a time window $T_{\rm a}-\tau_{\rm d},T_{\rm a}]$. If only the partial IE condition holds, there exists at least one inactive channel $\bm{\phi}_{{\rm s},j}$ for $j\in\{1,2,\cdots,N\}$, i.e., $\|\bm{\phi}_{{\rm s},j}\|\equiv 0$, and all elements in the $j$th row and the $j$th column of $\Psi(t)$ are 0. Fig. 2 illustrates the relationship between the channels $\bm{\phi}_{{\rm s},j}(t)$ and the excitation matrix $\Psi(t)$ with $N=7$. In this case, the full exciting strength $\sigma_{\min}(\Psi(t))\equiv 0$ [see the red dotted line in Fig. 1], such that online data need to be stored and updated by the excitation matrix $\Psi_{\zeta}$, which removes the effect that the current exciting strength $\sigma_{\min}(\Psi(t))$ is decided by inactive channels. Note that inactive channels can be determined by $\|\bm{\phi}_{{\rm s},j}\|<\mu$ with a small threshold $\mu\in\mathbb{R}^{+}$ instead of $\|\bm{\phi}_{{\rm s},j}\|=0$ in practice, as measurement noise may exist such that it is impossible to have $\|\bm{\phi}_{{\rm s},j}\|\neq 0$ even for inactive channels.
Existing composite learning methods only consider the IE case and make use of the excitation matrix $\Psi(t_{\rm e})$ in (14) with the auxiliary variable $\bm{q}(t_{\rm e})$ in (15) to compute the generalized prediction error $\bm{\xi}(t)$ in (19) directly [41]. However, $\Psi(t_{\rm e})$ can be discontinuous at some moments even the excitation strength $\sigma_{\rm c}(t)$ is continuous, which makes it impossible to directly obtain the high-order time derivatives of $\hat{\bm{\theta}}$, and the control law (6) may generate a discontinuous signal. Differently in the proposed composite learning HOT (23), $\Psi(t_{\rm e})$ in (14) is embedded in the stable filter $H(s)$ to generate the filtered excitation matrix $Q(t,t_{\rm e})$ in (17) and to construct $\bm{\xi}(t)$ in (19). In this manner, the high-order time derivatives $\hat{\bm{\theta}}^{(k+1)}$ in (24) and the control law (6) can be implemented continuously.
V Theoretical Guarantees
V-A Parameter Convergence Results
Let $\Phi_{{\rm f},\zeta}$ :=:= $\bm{\phi}_{{\rm f},k_{1}},\bm{\phi}_{{\rm f},k_{2}},\cdots,\bm{\phi}_{{\rm f},k_{N_{\zeta}}}]^{T}$ $\in\mathbb{R}^{N_{\zeta}\times n}$ denote an active sub-regressor of $\Phi_{\rm f}$ in (21), where $\bm{\phi}_{{\rm f},k_{j}}$ is the $k_{j}$th channel. Define a parameter estimation error $\tilde{\bm{\theta}}_{\zeta}$ :=:= $\tilde{\theta}_{k_{1}},\tilde{\theta}_{k_{2}},\cdots$, $\tilde{\theta}_{k_{N_{\zeta}}}]^{T}$ $\in\mathbb{R}^{N_{\zeta}}$ regarding active channels. The following theorem shows parameter convergence results of this study.
Theorem 1: Let $0,t_{\rm f})$ with $t_{\rm f}\in\mathbb{R}^{+}$ be the maximal time interval for the existence of solutions of the system (4). For any given unknown parameter vector $\bm{\theta}\in\Omega_{{\rm c}_{\theta}}$ and $\Gamma=\Gamma^{T}>0$, the composite learning law of $\hat{\bm{\theta}}$ in (23) ensures:
  1. 1.
    The estimation error $\tilde{\bm{\theta}}(t)$ is of $L_{\infty}$, $\forall t\geq 0$, and the general prediction error $\bm{\epsilon}(t)$ in (22) is of $L_{2}\cap L_{\infty}$, $\forall t\in[0,t_{\rm f})$;
  2. 2.
    The partial estimation error $\tilde{\bm{\theta}}_{\zeta}(t)$ exponentially converges to $\bm{0}$ if $t_{\rm f}\rightarrow\infty$, partial IE exists for constants $\sigma$, $T_{\rm a}\in\mathbb{R}^{+}$, and the index set $\mathcal{I}$ no longer changes on $t\in(T_{\rm a},\infty)$;
  3. 3.
    The estimation error $\tilde{\bm{\theta}}(t)$ exponentially converges to $\bm{0}$ if $t_{\rm f}\rightarrow\infty$ and IE exists for constants $\sigma$, $T_{\rm e}\in\mathbb{R}^{+}$.
The proof of Theorem 1 is given in Appendix A.
V-B Closed-Loop Stability Results
We introduce the following lemmas to facilitate the analysis of closed-loop stability, where Lemma 2 can be obtained by a simple extension of [13, Lemma 4].
Lemma 1 (local existence of solutions) [87]: Consider the system (4) under Assumptions 1–2 with the control law (6). For any given $\bm{x}(0)\in\Omega_{{\rm c}_{0}}\subset\mathbb{R}^{n}$ with $c_{0}\in\mathbb{R}^{+}$, there exists $c_{x}>c_{0}\in\mathbb{R}^{+}$, such that $\bm{x}(t)\in\Omega_{{\rm c}_{x}}\subset\mathbb{R}^{n}$, $\forall t\in[0,t_{\rm f})$.
Lemma 2 (convergence of stable filters) [13]: Consider $H(s)$ in (16) on $t\in[0,t_{\rm f})$ and $\Phi_{\rm f}$ in (20). For any given small $\delta\in\mathbb{R}^{+}$, there exists a sufficiently large $\alpha\in\mathbb{R}^{+}$ and $\alpha_{i}>\alpha$ with $i=1$ to n1n-1, so that $\|\Phi-\Phi_{\rm f}\|\leq\delta$, $\forall t\in[0,t_{\rm f})$.
The following theorem is established to show the stability of the closed-loop system (7) with (23).
Theorem 2: For the system (4) under Assumptions 1–3 driven by the CLBC law (6) and (23) with $\bm{x}(0)\in\Omega_{{\rm c}_{0}}$ and $\hat{\bm{\theta}}(0)\in\Omega_{{\rm c}_{\theta}}$, there exist proper control parameters $k_{{\rm c}1}$ to $k_{{\rm c}n}$ in (5) and filtered parameters $\alpha_{1}$ to $\alpha_{n-1}$ in (16), such that the equilibrium point $(\bm{e}$, $\tilde{\bm{\theta}})=\bm{0}$ of the closed-loop system (7) with (23) has:
  1. 1.
    Uniform ultimate boundedness (UUB) on $t$ $\in$ $0,\infty)$;
  2. 2.
    Partial exponential stability on $t\in[T_{\rm a},\infty)$ if partial IE in Definition 3 exists for constants $T_{\rm a}$, $\sigma$ $\in$ $\mathbb{R}^{+}$ and the set $\mathcal{I}$ no longer changes, where the tracking error $\bm{e}$ and the partial estimation error $\tilde{\bm{\theta}}_{\zeta}$ exponentially converge to $\bm{0}$;
  3. 3.
    Exponential stability on $t$ $\in$ $T_{\rm e},\infty)$ if IE in Definition 2 exists for constants $T_{\rm e}$, $\sigma\in\mathbb{R}^{+}$, where the tracking error $\bm{e}$ and the estimation error $\tilde{\bm{\theta}}$ exponentially converge to $\bm{0}$.
The proof of Theorem 2 is given in Appendix B.
Remark 3: In existing modular backstepping control methods, the nonlinear damping terms $k_{\rm{d}\it i}\|\bm{\psi}_{i}\|^{2}$ ($i=1$ to $n$) can ensure the boundedness of the system (4) in the absence of adaptation. In the proposed CLBC in (6) with (23), the non-update of some elements of the parameter estimate $\hat{\bm{\theta}}$ in (23) implies that the corresponding channels are inactive. In this case, partial IE in Definition 3 exists, and partial exponential stability of the closed-loop system without resorting to $k_{\rm{d}\it i}\|\bm{\psi}_{i}\|^{2}$ can be established under Algorithm 1. It is worth noting in Theorem 2 that if the condition “the index set $\mathcal{I}$ no longer changes on $t\in(T_{\rm a},\infty)$” is not met, partial exponential stability of the closed-loop system can still be established. Actually, the endless changes of some indices $k_{j}$ imply the establishment of partial PE.
Remark 4: The proposed CLBC in (6) with (23) has several connections and distinctions compared to the MRE-HOT in [43]. Similarly to the MRE-HOT, the swapping and the stable filter $H(s)$ in (16) are applied to generate the high-order time derivatives $\hat{\bm{\theta}}^{(k+1)}$ in (24) for counteracting the negative effect due to the transient of parameter estimation. The distinctions between the two controllers include:
  1. 1.
    The proposed CLBC introduces the filtered prediction error $\bm{\epsilon}$ in (22) to counteract the transient response caused by the modeling error $\Phi^{T}\tilde{\bm{\theta}}$ in (7) such that closed-loop stability is established without the nonlinear damping terms $k_{\rm{d}\it i}\|\bm{\psi}_{i}\|^{2}$ ($i=1$ to $n$), whereas the MRE-HOT resorts to $k_{\rm{d}\it i}\|\bm{\psi}_{i}\|^{2}$ to counteract $\Phi^{T}\tilde{\bm{\theta}}$ so as to establish UUB;
  2. 2.
    The proposed CLBC uses the extended regression equation (12) to obtain its generalized form (13), applies $H(s)$ to (13) to construct the generalized prediction error $\bm{\xi}$ in (19), and combines two prediction errors $\bm{\xi}$ and $\bm{\epsilon}$ for the composite learning HOT (23), whereas the MRE-HOT applies $H(s)$ directly to (12) to design an indirect adaptive law;
  3. 3.
    The proposed CLBC stores and forgets online data based on Algorithm 1 to ensure the monotonous non-decrease of the exciting strength $\sigma_{\rm c}$ in each excitation stage, enhancing exponential stability under partial IE or IE, but the MRE-HOT is hard to ensure the monotonous non-decrease of $\sigma_{\rm c}$, and its exponential stability relies on PE.
V-C Robustness Results
An external disturbance $\bm{d}(t)$ $\in$ $\mathbb{R}^{n}$ satisfying $\|\bm{d}(t)\|$ $\leq$ $\bar{d}$ is introduced to the system (4), where $\bar{d}$ $\in$ $\mathbb{R}^{+}$ is a constant. The closed-loop error system (7) can be rewritten into
$\dot{\bm{e}}=\Lambda\bm{e}+\Phi^{T}(\bm{x},\Theta_{n-1},{\bm{y}}_{{\rm r}n})\tilde{\bm{\theta}}+\bm{d}(t).$ (25)
Noting (10), consider the linear filtering operation
$\dot{\bm{d}}_{\rm s}(t)=\Lambda\bm{d}_{\rm s}(t)+\bm{d}(t)$ (26)
where $\bm{d}_{\rm s}(t)\in\mathbb{R}^{n}$ satisfies $\|\bm{d}_{\rm s}\|$ $\leq$ $\bar{d}/\sqrt{k_{\min}}$ with $k_{\min}$ :=:= $\min_{1\leq i\leq n}\{k_{{\rm c}i}\}$ $\in$ $\mathbb{R}^{+}$ as (26) is stable. Then, (11) becomes
$\bm{p}(t)=\Phi^{T}_{\rm s}(t)\bm{\theta}+\bm{d}_{\rm s}(t).$ (27)
Multiplying (27) by $\Phi_{\rm s}$, integrating the resultant equality at $\tau_{\rm d}, tt$], and using $\Psi$ in (14), one obtains
$\Psi(t)\bm{\theta}+\int_{t-\tau_{\rm d}}^{t}\Phi_{\rm s}(\tau)\bm{d}_{\rm s}(\tau)d\tau=\int_{t-\tau_{\rm d}}^{t}\Phi_{\rm s}(\tau)\bm{p}(\tau)d\tau.$ (28)
Applying the filter $H(s)$ in (16) to (28) and noting the generalized linearly parameterized model (17), one gets
$Q(t)\bm{\theta}+{\bm{d}}_{\rm g}(t)={\bm{q}}_{\rm f}(t)$ (29)
with ${\bm{d}}_{\rm g}(t)$ :=:= $H(s)[\int_{t-\tau_{\rm d}}^{t}\Phi_{\rm s}(\tau)\bm{d}_{\rm s}(\tau)d\tau]\in\mathbb{R}^{N}$. Let $\Omega_{{\rm c}_{x}}$ with ${\rm c}_{x}\in\mathbb{R}^{+}$ be a compact set of $\bm{x}$. Then, there exists a constant $\phi\in\mathbb{R}^{+}$ such that $\|\Phi_{\rm s}(t)\|\leq\phi$, $\forall\bm{x}\in\Omega_{{\rm c}_{x}}$. Noting $H(s)$ being stable with unit DC gain and combining the above results with $\|\bm{d}_{\rm s}\|$ $\leq$ $\bar{d}/\sqrt{k_{\min}}$, one gets $\|{\bm{d}}_{\rm g}(t)\|$ $\leq$ ${\bar{d}}_{\rm g}$, $\forall\bm{x}\in\Omega_{{\rm c}_{x}}$ with ${\bar{d}}_{\rm g}$ :=:= $\tau_{\rm d}\phi\bar{d}/\sqrt{k_{\min}}$. Also, applying $H(s)$ to the closed-loop error system (7) and noting (20), one gets
$\bm{z}(t)=\Phi_{\rm f}^{T}(t)\bm{\theta}+{\bm{d}}_{\rm f}(t)$ (30)
where ${\bm{d}}_{\rm f}$ :=:= $H(s)[\bm{d}]$ satisfies $\|\bm{d}_{\rm f}\|\leq\bar{d}$ due to the properties of $H(s)$. The following theorem is established to show the robustness results of the proposed CLBC law (6) and (23).
Theorem 3: For the system (4) with the additive disturbance $\bm{d}$ and Assumptions 1–3 driven by the CLBC law (6) and (23) with $\bm{x}(0)\in\Omega_{{\rm c}_{0}}$ and $\hat{\bm{\theta}}(0)\in\Omega_{{\rm c}_{\theta}}$, there exist suitable control parameters $k_{{\rm c}1}$ to $k_{{\rm c}n}$ in (5) and filtered parameters $\alpha_{1}$ to $\alpha_{n-1}$ in (16), such that the equilibrium point $(\bm{e}$, $\tilde{\bm{\theta}})=\bm{0}$ of the closed-loop system (25) with (23) has:
  1. 1.
    UUB stability on $t$ $\in$ $0,\infty)$;
  2. 2.
    Partial practical exponential stability on $t$ $\in$ $T_{\rm a},\infty)$ when partial IE exists for constants $T_{\rm a}$, $\sigma\in\mathbb{R}^{+}$ and the set $\mathcal{I}$ no longer changes, where the tracking error $\bm{e}$ and the partial estimation error $\tilde{\bm{\theta}}_{\zeta}$ exponentially converge to a small neighborhood of $\bm{0}$ dominated by $k_{{\rm c}i}$, $\kappa$, and $\bar{d}$;
  3. 3.
    Practical exponential stability on $t$ $\in$ $T_{\rm e},\infty)$ if IE exists for constants $T_{\rm e}$, $\sigma$ $\in$ $\mathbb{R}^{+}$, where the tracking error $\bm{e}$ and the estimation error $\tilde{\bm{\theta}}$ exponentially converge to a small neighborhood of $\bm{0}$ dominated by $k_{{\rm c}i}$, $\kappa$, and $\bar{d}$.
The proof of Theorem 3 is given in Appendix C.
VI Simulation Studies
VI-A Stability and Convergence Comparisons
This section is devoted to verifying the exponential stability and parameter convergence of the proposed CLBC in (6) with (23) under various excitation conditions. Consider a mass-spring-damping model as follows [2]:
$\left\{\begin{array}[]{l}\dot{x}_{1}=x_{2},\\ \dot{x}_{2}=x_{3}+\bm{\varphi}_{2}^{T}(\bm{x}_{2})\bm{\theta},\\ \dot{x}_{3}=u,\\ y=x_{1}\end{array}\right$.
where $x_{1}\in\mathbb{R}$ denotes a mass position, $u\in\mathbb{R}$ is a control input, and $\bm{\theta}\in\mathbb{R}^{3}$ is a unknown parameter vector. Noting (4), one has $\bm{\varphi}_{2}({\bm{x}}_{2})=[-x_{2},-x_{1},-x_{2}^{3}]^{T}$ and $\bm{\varphi}_{1}(x_{1})=\bm{\varphi}_{3}({\bm{x}})=\bm{0}$.
Set the control parameters $k_{{\rm c}i}=1$ with $i=1,2,3$, $\Gamma$ == $3I$, $\kappa=1$, $\tau_{\rm d}=3$, $\hat{\bm{\theta}}(0)=\bm{0}$, and $\sigma=10^{-4}$, and choose the stable filter $H(s)=25/(s^{2}+10s+25)$ in (16). Gaussian white noise with 0 mean and 0.001 standard deviation is added to the measurement of the states $x_{i}$. The MRE-HOT in [43] and the composite learning dynamic surface control (CL-DSC) in [49] are selected as baseline controllers, where we set the damping parameters $k_{\rm{d}\it i}=$ 0.1 for the MRE-HOT, the stable filter $L(s)$ == $20/(s+20)$ for the CL-DSC, and the other shared parameters to be the same values for fair comparisons.
Figure 3: Performance comparisons of three controllers for the tracking problem under the PE condition. (a) The absolute tracking errors $|e_{1}|$. (b) The estimation error norms $\|\tilde{\bm{\theta}}\|$. (c) The control inputs $u$. (d) The exciting strengths $\sigma_{\rm c}$.
Case 1: Tracking with PE. Consider a tracking problem under PE with the reference trajectory $y_{\rm r}=1.5\sin(0.5t)$, the desired parameter $\bm{\theta}=[0.1,0.5,1.5]^{T}$, and the initial state $\bm{x}(0)$ == $0.6,0,0]^{T}$. Performance comparisons of the three controllers are depicted in Figs. 3. It is observed that the estimation error $\tilde{\bm{\theta}}$ by the CL-DSC has a steady-state error since the time derivative of each virtual control input is estimated by the filter $L(s)$ [see Fig. 3(a)], $\tilde{\bm{\theta}}$ by the MRE-HOT converges to $\bm{0}$ after running 30 s [see Fig. 3(a)] as PE is fulfilled in this case, and $\tilde{\bm{\theta}}$ by the proposed CLBC exhibits the rapid convergence to $\bm{0}$ [see Fig. 3(a)], which validates its strong learning capability. Besides, the proposed CLBC exhibits much better tracking performance than the CL-DSC and MRE-HOT [see Fig. 3(b)] because the exciting strength $\sigma_{\rm c}$ of the CLBC is monotonic non-decreasing and keeps a high level throughout [see Fig. 3(d)]. Moreover, the control inputs $u$ by the proposed CLBC and the MRE-HOT are not affected by the measurement noise throughout and are comparable after their estimation errors $\tilde{\bm{\theta}}$ converge to $\bm{0}$ [see Fig. 3(c)], but $u$ by the CL-DSC is seriously polluted by the measurement noise [see Fig. 3(c)], which results in much worse tracking performance [see Fig. 3(b)].
Figure 4: Performance comparisons of three controllers for the regulation problem under partial IE or IE condition. (a) The partial estimation errors $\|\tilde{\bm{\theta}}_{\zeta}\|$. (b) The estimation errors $\|\tilde{\bm{\theta}}\|$. (c) The absolute tracking error norms $|e_{1}|$. (d) The control inputs $u$. (e) The exciting strengths $\sigma_{\rm c}$.
Case 2: Regulation with partial IE or IE. Consider a regulation problem under the partial IE or IE condition with the desired parameter $\bm{\theta}=[0.4,0.5,0.1]^{T}$ and the initial state $\bm{x}(0)=\bm{0}$, where the reference trajectories $y_{\rm r}$, $\dot{y}_{\rm r}$, $\ddot{y}_{\rm r}$, and $\dddot{y}_{\rm r}$ are generated by a reference model $y_{\rm r}(t)$ == $\frac{a_{0}}{a(s)}[r(t)]$ with $a_{0}=16$, $a(s)=s^{4}+8s^{3}+24s^{2}+32s+16$, and
[…]
such that partial IE exists in $t$ $\in$ [0,60) s, and IE exists in $t$ $\in$ [60,120] s. Performance comparisons of the three controllers are depicted in Fig. 4 with $\tilde{\bm{\theta}}_{\zeta}$ :=:= $\tilde{\theta}_{1},\tilde{\theta}_{2}]^{T}$. It is observed that the CL-DSC exhibits the rapid convergence of the estimation error $\tilde{\bm{\theta}}$ at $t\in[60,120]$ s due to the establishment of IE but still has a steady-state error [see Fig. 4(b)], the MRE-HOT does not show the convergence of $\tilde{\bm{\theta}}$ to $\bm{0}$ after 60 s [see Fig. 4(b)] due to the lack of PE, and the proposed CLBC shows the convergence of partial elements $\tilde{\theta}_{1}$ and $\tilde{\theta}_{2}$ to 0 at $t\in[0,60)$ s and then all elements $\tilde{\theta}_{i}$ ($i=1$ to $3$) to 0 after $t=60$ s [see Figs. 4(a) and (b)], which is consistent with the theoretical result in Theorem 1. Regarding the tracking performance, the proposed CLBC owns the highest tracking accuracy after $\tilde{\bm{\theta}}$ converges to $\bm{0}$ [see Fig. 4(c)]. Moreover, the control inputs $u$ by the proposed CLBC and the MRE-HOT are comparable in this case, and the control input $u$ by the CL-DSC is sensitive to the measurement noise [see Fig. 4(d)]. It is worth noting that the exciting strengths $\sigma_{\rm c}$ by the CL-DSC and MRE-HOT are 0 at $t\in[0,60)$ s due to the inactive channel $\bm{\phi}_{3}$ [see Fig. 4(e)], but their different control designs lead in obviously different tracking results as shown above.
VI-B Transient Performance Comparisons
This section is devoted to demonstrating that the transient performance of the proposed CLBC in (6) with (23) can still be guaranteed without the nonlinear damping terms $k_{\rm{d}\it i}\|\bm{\psi}_{i}\|^{2}$ ($i=1$ to $n$). Consider a second-order system [84]
$\left\{\begin{array}[]{l}\dot{x}_{1}=x_{2}+\varphi_{1}(x_{1})\theta,\\ \dot{x}_{2}=u,\\ y=x_{1}\end{array}\right$.
with $u\in\mathbb{R}$, $\varphi_{1}(x_{1})=x_{1}^{2}$, $\theta=2$, and $\bm{x}(0)=[0.6,0]^{T}$. Note that for existing modular backstepping control methods, if the damping coefficients $k_{\rm{d}\it i}$ are close to 0, the transient performance of the above system can be deteriorated by the modeling error term $\Phi^{T}\tilde{\theta}$ regardless of the convergence of the estimation error $\tilde{\theta}$ [84]. The reference trajectories $y_{\rm r}$, $\dot{y}_{\rm r}$, and $\ddot{y}_{\rm r}$ are generated by $y_{\rm r}(t)$ == $\frac{a_{0}}{a(s)}[r(t)]$ with $a_{0}=1$, $a(s)=s^{3}+3s^{2}+3s+1$, and $r(t)=\sin(2t)$. Set the control parameters $k_{{\rm c}1}=k_{{\rm c}2}=1$, $\Gamma=1$, $\kappa=1$, $\tau_{\rm d}=1$, $\hat{\theta}(0)=0$, and $\sigma=10^{-4}$, and choose the stable filter $H(s)=25/(s+25)$. The HOT adaptive backstepping control (HOT-ABC) in [30] and the MRE-HOT in [43] are selected as baseline controllers, where their shared parameters are set to be the same for fair comparisons.
[…]
Performance comparisons of transient tracking and parameter convergence of the three controllers are exhibited in Fig. 5, where both the damping parameters $k_{\rm{d}\it 1}$ and $k_{\rm{d}\it 2}$ increase from $0.01$ to $0.19$ with a step size of $0.03$, and the arrows indicate the increasing direction $k_{\rm{d}\it 1}$ and $k_{\rm{d}\it 2}$. It is observed that the transient results of the HOT-ABC and MRE-HOT are deteriorated by decreasing $k_{\rm{d}\it 1}$ and $k_{\rm{d}\it 2}$ [see green and red dash lines in Fig. 5(a)]. For the MRE-HOT, increasing $k_{\rm{d}\it 1}$ and $k_{\rm{d}\it 2}$ reduces the initial oscillations of the tracking error $|e_{1}|$ [see green dash lines in Fig. 5(a)] but also slows down parameter convergence [see green dash lines in Fig. 5(b)]. The proposed CLBC provides the convergence of $|e_{1}|$ to 0 with the smallest initial oscillations [see Fig. 5(a)] and achieves the fastest parameter convergence [see Fig. 5(b)], which implies that: 1) the transient tracking and parameter convergence of the HOT-ABC and MRE-HOT are sensitive to damping parameters; 2) the transient performance of the proposed CLBC is enhanced by introducing the extra prediction error $\bm{\epsilon}$ in (22) without resorting to nonlinear damping terms; 3) The combination of two prediction errors $\bm{\xi}$ in (19) and $\bm{\epsilon}$ in (22) accelerates parameter convergence.
VII Conclusions
This paper has presented a feasible modular backstepping control strategy named CLBC for strict-feedback uncertain nonlinear systems. The proposed composite learning HOT allows the exact implementation of the high-order time derivatives of parameter estimates and the offset of modeling errors, such that the transient performance can be guaranteed without resorting to nonlinear damping terms or high control gains. The proposed algorithm of staged exciting strength maximization ensures that the exciting strength is monotonously non-decreasing in each excitation stage, such that the exponential stability of the closed-loop system with parameter convergence is obtainable under the much weaker condition of IE or partial IE. Simulation studies have validated that the proposed CLBC greatly outperforms two state-of-the-art modular backstepping controllers, namely HOT-ABC and MRE-HOT, in both parameter estimation and control. Further work would focus on nonlinear parameterization problems and robot control based on the proposed method.
Appendix A The proof of Theorem 1
Proof.
1) Since one has $\bm{\xi}=Q(t,t_{\rm e})\tilde{\bm{\theta}}$, $\bm{\epsilon}=\Phi_{\rm f}^{T}\tilde{\bm{\theta}}$, and $\dot{\tilde{\bm{\theta}}}=-\dot{\hat{\bm{\theta}}}$, the composite learning law (23) becomes
[…]
Choose a Lyapunov function candidate
$V_{\theta}=\tilde{\bm{\theta}}^{T}\Gamma^{-1}\tilde{\bm{\theta}}$. (32)
Differentiating $V_{\theta}$ regarding $t$ and using (31) yields
$\dot{V}_{\theta}\leq-2\kappa\tilde{\bm{\theta}}^{T}Q(t,t_{\rm e})\tilde{\bm{\theta}}-2\tilde{\bm{\theta}}^{T}\Phi_{\rm f}\Phi_{\rm f}^{T}\tilde{\bm{\theta}}$.
Because $Q(t,t_{\rm e})$ is positive semidefinite, i.e., $Q(t,t_{\rm e})\geq 0$, one has $\tilde{\bm{\theta}}^{T}Q(t,t_{\rm e})\tilde{\bm{\theta}}\geq 0$ such that
$\dot{V}_{\theta}\leq-2\tilde{\bm{\theta}}^{T}\Phi_{\rm f}\Phi_{\rm f}^{T}\tilde{\bm{\theta}},\forall t\in[0,t_{\rm f}).$ (33)
Noting $V_{\theta}(t)\geq 0$ and $\dot{V}_{\theta}(t)\leq 0$ from (33), one obtains $0\leq V_{\theta}(t)\leq V_{\theta}(0)$, $\forall t\geq 0$. Integrating each side of (33) over $[0,t]$ and using $\bm{\epsilon}=\Phi_{\rm f}^{T}\tilde{\bm{\theta}}$, one obtains
$\int_{0}^{t}\bm{\epsilon}^{2}(\tau)d\tau\leq(V_{\theta}(0)-V_{\theta}(t))/2.$
Noting $V_{\theta}(t)\in L_{\infty}$ and $0\leq V_{\theta}(t)\leq V_{\theta}(0)$, the above result implies $\bm{\epsilon}\in L_{2}$, $\forall t\in[0,t_{\rm f})$. Using $V_{\theta}(t)\leq V_{\theta}(0)$ yields
$\tilde{\bm{\theta}}^{T}(t)\Gamma^{-1}\tilde{\bm{\theta}}(t)\leq\tilde{\bm{\theta}}^{T}(0)\Gamma^{-1}\tilde{\bm{\theta}}(0).$
After some algebraic operations, the above result leads to
$\|\tilde{\bm{\theta}}(t)\|\leq\sqrt{\lambda_{\max}(\Gamma)/\lambda_{\min}(\Gamma)}\|\tilde{\bm{\theta}}(0)\|$
implying $\tilde{\bm{\theta}}(t)\in L_{\infty}$. Since the solutions of (4) only exist on $t\in[0,t_{\rm f})$, one obtains $\bm{x}(t)$ and $\hat{\bm{\theta}}^{(k)}(t)\in L_{\infty}$ with $k=1$ to n1n-1, $\forall t\in[0,t_{\rm f})$, implying $\Phi_{\rm f}\in L_{\infty}$, $\forall t\in[0,t_{\rm f})$. It follows from $\bm{\epsilon}=\Phi_{\rm f}^{T}\tilde{\bm{\theta}}$ that $\bm{\epsilon}\in L_{\infty}$, $\forall t\in[0,t_{\rm f})$. In summary, one has $\tilde{\bm{\theta}}(t)\in L_{\infty}$, $\forall t\geq 0$ and $\bm{\epsilon}(t)\in L_{2}\cap L_{\infty}$, $\forall t\in[0,t_{\rm f})$.
2) Consider the parameter convergence problem under partial IE and $t_{\rm f}\rightarrow\infty$. The composite learning law (23) with respect to active channels is given by
$\dot{\tilde{\bm{\theta}}}_{\zeta}=-\Gamma_{\zeta}\left(\Phi_{{\rm f},\zeta}\Phi_{{\rm f},\zeta}^{T}\tilde{\bm{\theta}}_{\zeta}+\kappa Q_{\zeta}(t,t_{\rm e}){\tilde{\bm{\theta}}}_{\zeta}\right),t\geq T_{\rm a}$ (34)
with $Q_{\zeta}(t,t_{\rm e})$ :=:= $H(s)[\Psi_{\zeta}(t_{\rm e})]$, in which $\Gamma_{\zeta}\in\mathbb{R}^{N_{\zeta}\times N_{\zeta}}$ is a positive-definite diagonal matrix of learning rates with respect to active channels. Let a Lyapunov function candidate be
$V_{\theta,\zeta}=\tilde{\bm{\theta}}_{\zeta}^{T}\Gamma^{-1}_{\zeta}\tilde{\bm{\theta}}_{\zeta}.$ (35)
Differentiating $V_{\theta,\zeta}$ with respect to $t$ and applying (34) to the resulting expression, one obtains
${\dot{V}}_{\theta,\zeta}\leq$ $-2\tilde{\bm{\theta}}_{\zeta}^{T}\Phi_{{\rm f},\zeta}\Phi_{{\rm f},\zeta}^{T}\tilde{\bm{\theta}}_{\zeta}-2\kappa\tilde{\bm{\theta}}_{\zeta}^{T}Q_{\zeta}(t,t_{\rm e}){\tilde{\bm{\theta}}}_{\zeta}$
$\leq$ $-2\kappa\tilde{\bm{\theta}}_{\zeta}^{T}Q_{\zeta}(t,t_{\rm e}){\tilde{\bm{\theta}}}_{\zeta}.$
Since partial IE exists for the constant $\sigma$, one gets $\Psi_{\zeta}(t_{\rm e})\geq\sigma_{\rm c}(T_{\rm a})I\geq\sigma I$ from (18). As $H(s)$ is stable with unit DC gain, there exists a constant $\sigma^{*}$ $\in$ $\mathbb{R}^{+}$ with $\sigma^{*}\leq\sigma$ such that
$Q_{\zeta}(t,t_{\rm e})=H(s)[\Psi_{\zeta}(t_{\rm e})]\geq\sigma^{*}I.$
It follows from the above two results that
$\dot{V}_{\theta,\zeta}\leq-2\kappa\sigma^{*}\tilde{\bm{\theta}}_{\zeta}^{T}{\tilde{\bm{\theta}}}_{\zeta}\leq-2\kappa\sigma^{*}\lambda_{\min}(\Gamma_{\zeta})\tilde{\bm{\theta}}_{\zeta}^{T}\Gamma^{-1}_{\zeta}{\tilde{\bm{\theta}}}_{\zeta}.$
Thus, one immediately obtains
$\dot{V}_{\theta,\zeta}(t)\leq-k_{\zeta}V_{\theta,\zeta}(t),t\geq T_{\rm a}$
with $k_{\zeta}$ :=:= $2\kappa\sigma^{*}\lambda_{\min}(\Gamma_{\zeta})\in\mathbb{R}^{+}$. Applying [88, Lemma A.3.2] to solve the above inequality, one obtains
$V_{\theta,\zeta}(t)\leq V_{\theta,\zeta}(T_{\rm a})e^{-k_{\zeta}(t-T_{\rm a})}.$
Combining (35) with the above result yields
$\|\tilde{\bm{\theta}}_{\zeta}(t)\|^{2}\leq$ $\lambda_{\max}(\Gamma_{\zeta})\tilde{\bm{\theta}}_{\zeta}^{T}(t)\Gamma_{\zeta}^{-1}\tilde{\bm{\theta}}_{\zeta}(t)=\lambda_{\max}(\Gamma_{\zeta})V_{\theta,\zeta}(t)$
$\leq$ $\lambda_{\max}(\Gamma_{\zeta})V_{\theta,\zeta}(T_{\rm a})e^{-k_{\zeta}(t-T_{\rm a})},\forall t\geq T_{\rm a}$
which implies that the partial estimation error $\tilde{\bm{\theta}}_{\zeta}(t)$ exponentially converges to $\bm{0}$ on $t\in[T_{\rm a},\infty)$.
3) The proof of parameter convergence under IE is similar to that under partial IE, so we omit some similar steps. Applying $V_{\theta}$ in (32) and the IE condition $\Psi(T_{\rm e})\geq\sigma I$, one obtains
$\|\tilde{\bm{\theta}}(t)\|^{2}\leq$ $\lambda_{\max}(\Gamma)\tilde{\bm{\theta}}^{T}\Gamma^{-1}\tilde{\bm{\theta}}=\lambda_{\max}(\Gamma)V_{\theta}(t)$
$\leq$ $\lambda_{\max}(\Gamma)V_{\theta}(T_{\rm e})e^{-k_{\theta}(t-T_{\rm e})},\forall t\geq T_{\rm e}$
where $k_{\theta}$ :=:= $2\kappa\sigma^{*}\lambda_{\min}(\Gamma)$, and $\sigma^{*}\in\mathbb{R}^{+}$ with $\sigma^{*}\leq\sigma$ satisfies
$Q(t,t_{\rm e})=H(s)[\Psi(t_{\rm e})]\geq H(s)[\Psi(T_{\rm e})]\geq\sigma^{*}I$ (36)
which implies that the parameter estimation error $\tilde{\bm{\theta}}(t)$ exponentially converge to $\bm{0}$ on $t\in[T_{\rm e},\infty)$. ∎
Appendix B The proof of Theorem 2
Proof.
1) Choose a Lyapunov function candidate
$V(\bm{w})=\bm{e}^{T}\bm{e}/2+(1+p)\tilde{\bm{\theta}}^{T}\Gamma^{-1}\tilde{\bm{\theta}}/2$ (37)
with $\bm{w}$ :=:= $\bm{e}^{T},\tilde{\bm{\theta}}^{T}]^{T}\in\mathbb{R}^{n+N}$, where $p\in\mathbb{R}^{+}$ is a constant. There exist constants $\lambda_{\rm a}$ :=:= $\min\{1/2,(1+p)/(2\lambda_{\max}(\Gamma))\}$ and $\lambda_{\rm b}$ :=:= $\min\{1/2,(1+p)/(2\lambda_{\min}(\Gamma))\}$ such that
$\lambda_{\rm a}\|\bm{w}\|^{2}\leq V(\bm{w})\leq\lambda_{\rm b}\|\bm{w}\|^{2}.$ (38)
Noting (7) and (25), one obtains the closed-loop system
[…]
Differentiating $V$ with respect to $t$ yields
$\dot{V}=(\dot{\bm{e}}^{T}\bm{e}+\bm{e}^{T}\dot{\bm{e}})/2+(1+p)\tilde{\bm{\theta}}^{T}\Gamma^{-1}\dot{\tilde{\bm{\theta}}}$.
Applying (41) to the above formula, one obtains
$\dot{V}=$ $\bm{e}^{T}(-K\bm{e}+\Phi^{T}\tilde{\bm{\theta}})-(1+p)\tilde{\bm{\theta}}^{T}(\Phi_{{\rm f}}\Phi_{{\rm f}}^{T}+\kappa Q(t,t_{\rm e}))\tilde{\bm{\theta}}$ (42)
with $K:=$ diag($k_{{\rm c}1},k_{{\rm c}2},\cdots,k_{{\rm c}n}$), which can be rewritten into
$\dot{V}=$ $-\bm{e}^{T}(K-I/4)\bm{e}-\kappa(1+p)\tilde{\bm{\theta}}^{T}Q(t,t_{\rm e}){\tilde{\bm{\theta}}}$
$+\bm{e}^{T}(\Phi-\Phi_{\rm f})\tilde{\bm{\theta}}-\bm{e}^{T}\bm{e}/4$
$-\tilde{\bm{\theta}}^{T}\Phi_{\rm f}\Phi_{\rm f}^{T}\tilde{\bm{\theta}}+\bm{e}^{T}\Phi_{\rm f}^{T}\tilde{\bm{\theta}}-p\tilde{\bm{\theta}}^{T}\Phi_{\rm f}\Phi_{\rm f}^{T}\tilde{\bm{\theta}}$
where the second line of the above formula satisfies
$-\bm{e}^{T}\bm{e}/4-\tilde{\bm{\theta}}^{T}\Phi_{\rm f}\Phi_{\rm f}^{T}\tilde{\bm{\theta}}+\bm{e}^{T}\Phi_{\rm f}^{T}\tilde{\bm{\theta}}\leq-\|\bm{e}/2-\Phi_{\rm f}^{T}\tilde{\bm{\theta}}\|^{2}\leq 0.$
Then, it is straightforward to get
$\dot{V}\leq$ $-\bm{e}^{T}(K-I/4)\bm{e}-\kappa(1+p)\tilde{\bm{\theta}}^{T}Q(t,t_{\rm e}){\tilde{\bm{\theta}}}$
$+\bm{e}^{T}(\Phi-\Phi_{\rm f})\tilde{\bm{\theta}}$. (43)
From Lemma 2, one gets $\|\Phi-\Phi_{\rm f}\|\leq\delta$, $\forall t\in[0,t_{\rm f})$, in which $t_{\rm f}\in\mathbb{R}^{+}$ is the moment that $\bm{x}(t)$ leaves $\Omega_{{\rm c}_{x}}$ for the first time. According to Theorem 1, one gets $\tilde{\bm{\theta}}(t)\in L_{\infty}$, $\forall t\in[0,\infty)$, such that there exists a constant $c_{\theta}\in\mathbb{R}^{+}$ that satisfies $\|\tilde{\bm{\theta}}(t)\|\leq c_{\theta}$, $\forall t\in[0,\infty)$. Applying these results to the above inequality and letting $k_{{\rm c}i}>1/4$, one obtains
$\dot{V}\leq$ $-k_{\rm c}\|\bm{e}\|^{2}+\delta c_{\theta}\|\bm{e}\|$
with $k_{\rm c}$ :=:= $\min_{1\leq i\leq n}\{k_{{\rm c}i}\}-1/4\in\mathbb{R}^{+}$. Applying Young’s inequality $ab\leq a^{2}/2+b^{2}/2$ with $a=\sqrt{k_{\rm c}}\|\bm{e}\|$ and $b=\delta c_{\theta}/\sqrt{k_{\rm c}}$ to the above inequality, one obtains
$\dot{V}\leq-k_{\rm c}\|\bm{e}\|^{2}/2+(\delta c_{\theta})^{2}/(2k_{\rm c})$ (44)
which is valid on $\bm{x}\in\Omega_{{\rm c}_{x}}$ for $t\in[0,t_{\rm f})$. It is observed from (B) and (44) that the prediction error $\bm{\epsilon}$ in (22) counteracts the modeling error term $\Phi^{T}\tilde{\bm{\theta}}$, leading to (44). Thus, $\bm{e}(t)$ converges to a steady-state bound subject to $\delta$ on $t\in[0,t_{\rm f})$. Based on (37) and $\|\tilde{\bm{\theta}}(t)\|\leq c_{\theta}$, $\forall t\in[0,\infty)$, (44) is rewritten into
$\dot{V}(t)\leq$ $-k_{\rm c}\|\bm{e}\|^{2}/2-k_{\rm c}(1+p)\tilde{\bm{\theta}}^{T}\Gamma^{-1}\tilde{\bm{\theta}}/2$
$+k_{\rm c}(1+p)c_{\theta}^{2}/(2\lambda_{\min}(\Gamma))+(\delta c_{\theta})^{2}/(2k_{\rm c})$
$=$ $-k_{\rm c}V(t)+k_{\rm c}(1+p)c_{\theta}^{2}/(2\lambda_{\min}(\Gamma))+(\delta c_{\theta})^{2}/(2k_{\rm c})$
$=$ $-k_{\rm c}V(t)/2-k_{\rm c}(V(t)-\eta(k_{\rm c},\alpha,\Gamma))/2$
with $\eta(k_{\rm c},\alpha,\Gamma)$ :=:= $(1+p)c_{\theta}^{2}/\lambda_{\min}(\Gamma)+(\delta c_{\theta}/k_{\rm c})^{2}\in\mathbb{R}^{+}$. Let $\Omega_{r}$ :=:= $\{{\bm{y}}_{{\rm r}n}|\dot{y}_{\rm r}$, $\cdots$, $y^{(n-1)}_{\rm r}\in\Omega_{{\rm c}_{\rm r}}\}\subset\mathbb{R}^{n}$, $\Omega_{{\rm c}_{w}}$ :=:= $\Omega_{{\rm c}_{x}}\cap\Omega_{r}\times\Omega_{{\rm c}_{\theta}}$, and $\Omega_{{\rm c}_{w0}}$ :=:= $\Omega_{{\rm c}_{0}}\cap\Omega_{r}\times\Omega_{{\rm c}_{\theta}}$ such that $\Omega_{{\rm c}_{w0}}\subset\Omega_{{\rm c}_{w}}$. It is implied from the above inequality that
$\dot{V}(t)\leq-k_{\rm c}V(t)/2,\forall V(t)\geq\eta$ (45)
on $\bm{w}(t)\in\Omega_{{\rm c}_{w}}$ and $t\in[0,t_{\rm f})$. Based on the results presented in (38) and (45), the UUB Theorem [90, Th. 4.5] is applied to conclude that if $\eta<\sqrt{c_{w0}/\lambda_{\rm b}}$, then the closed-loop system (41) achieves UUB stability, where $\bm{e}(t)$, $\tilde{\bm{\theta}}(t)\in$ $L_{\infty}$, $\forall t\geq 0$ implying $t_{\rm f}=\infty$ so that the solution $\bm{w}(t)$ is unique, $\forall t\geq 0$ [90, Lemma 3.1]. It is clear from the definition of $\eta$ that given any $c_{0}\in\mathbb{R}^{+}$, there certainly exist suitable $k_{\rm c}$, $\alpha$, and $\Gamma$ such that $\eta<\sqrt{c_{w0}/\lambda_{\rm b}}$ holds. Besides, the transient bound for $\bm{e}$ derived by [90, Th. 4.5] is given as follows:
$\|\bm{e}(t)\|\leq$ $\sqrt{\lambda_{\rm b}/\lambda_{\rm a}}\max\{(\|\bm{e}(0)\|+c_{\theta})e^{-(k_{\rm c}\lambda_{\rm a})/(4\lambda_{\rm b})t},$
$\sqrt{\eta/\lambda_{\rm b}}\},\forall t\geq 0.$ (46)
Using Assumptions 1–2 and $\bm{e}(t)$, $\tilde{\bm{\theta}}(t)\in$ $L_{\infty}$, $\forall t\geq 0$, one gets $\bm{x}(t)$ $\in$ $L_{\infty}$, $\forall t\geq 0$ in (7). As the signals $\Phi_{\rm f}$, $\bm{z}$, $Q$, and $\bm{q}_{\rm f}$ in (23) are bounded, and are filtered by $H(s)$ in (16), their time derivatives up to the (n2n-2)th order are also bounded according to the input-to-state stability property of the stable filter $H(s)$ [87, 43]. As the high-order time derivatives of $\hat{\bm{\theta}}$ in (24) are computed as a weighted sum of these signals and their high-order time derivatives, one concludes ${\hat{\bm{\theta}}}^{(k)}\in L_{\infty}$ for all relevant $k$. Thus, the equilibrium point $(\bm{e}$, $\tilde{\bm{\theta}})=\bm{0}$ of the closed-loop system (7) with (23) is UUB stable on $t\in[0,\infty)$.
2) Consider the control problem under partial IE on $t$ $\in$ $T_{\rm a}$, $\infty)$. If only partial IE in Definition 3 exists for some constants $\sigma$, $T_{\rm a}\in\mathbb{R}^{+}$, and the index set $\mathcal{I}$ no longer changes on $t\in[T_{\rm a},\infty)$, there exist some inactive channels $\bm{\phi}_{j}(t)$ such that $\|\bm{\phi}_{j}(t)\|\equiv 0$, $\forall t\in[T_{\rm a},\infty)$. Let $\Phi_{\zeta}$ :=:= $\bm{\phi}_{1},\bm{\phi}_{2},\cdots,\bm{\phi}_{{N_{\zeta}}}]^{T}$ $\in\mathbb{R}^{N_{\zeta}\times n}$ be an active sub-regressor of $\Phi$ with $\bm{\phi}_{j}$ being the $j$th column of $\Phi^{T}$, $\Phi_{{\rm f},\zeta}$ :=:= $\bm{\phi}_{{\rm f},1},\bm{\phi}_{{\rm f},2},\cdots,\bm{\phi}_{{\rm f},{N_{\zeta}}}]^{T}$ $\in\mathbb{R}^{N_{\zeta}\times n}$ denote an active sub-regressor of $\Phi_{\rm f}$ with $\bm{\phi}_{{\rm f},j}$ being the $j$th column of $\Phi_{\rm f}^{T}$, and $\tilde{\bm{\theta}}_{\zeta}$ :=:= $\tilde{\theta}_{1},\tilde{\theta}_{2},\cdots,\tilde{\theta}_{N_{\zeta}}]^{T}\in\mathbb{R}^{N_{\zeta}}$ denote an estimation error regarding active channels. Thus, the regressors $\Phi$ and $\Phi_{\rm f}$ can be represented by sub-regressors $\Phi_{\zeta}$ and $\Phi_{{\rm f},\zeta}$, respectively, i.e.,
$\Phi_{\rm f}=[\Phi_{{\rm f},\zeta}^{T},\underbrace{\bm{0},\bm{0},\cdots,\bm{0}}_{N-N_{\zeta}}]^{T},\Phi=[\Phi_{\zeta}^{T},\underbrace{\bm{0},\bm{0},\cdots,\bm{0}}_{N-N_{\zeta}}]^{T},$
and the closed-loop system (41) can be rewritten into
$\left\{\begin{array}[]{l}\dot{\bm{e}}=\Lambda\bm{e}+\Phi_{\zeta}^{T}\tilde{\bm{\theta}}_{\zeta}\\ \dot{\tilde{\bm{\theta}}}_{\zeta}=-\Gamma_{\zeta}\left(\Phi_{{\rm f},\zeta}\Phi_{{\rm f},\zeta}^{T}\tilde{\bm{\theta}}_{\zeta}+\kappa Q_{\zeta}(t,t_{\rm e}){\tilde{\bm{\theta}}}_{\zeta}\right)\\ \dot{\tilde{\bm{\theta}}}_{0}=\bm{0}\in\mathbb{R}^{N-N_{\zeta}}\end{array}\right$. (50)
where $\tilde{\bm{\theta}}_{0}$ :=:= $\tilde{\theta}_{N_{\zeta}+1},\tilde{\theta}_{N_{\zeta}+2},\cdots,\tilde{\theta}_{N}]^{T}\in\mathbb{R}^{N-N_{\zeta}}$ is an estimation error on inactive channels, and $Q_{\zeta}(t,t_{\rm e})$ is defined in (28). Note that the estimation error $\tilde{\bm{\theta}}_{0}(t)\equiv\tilde{\bm{\theta}}_{0}(T_{\rm a})$, $\forall t\geq T_{\rm a}$, and the tracking error $\bm{e}$ and the estimation error $\tilde{\bm{\theta}}_{\zeta}$ are not affected by $\tilde{\bm{\theta}}_{0}$. Thus, choose a new Lyapunov function candidate
$V_{\zeta}=\frac{1}{2}\bm{e}^{T}\bm{e}+\frac{1}{2}(1+p)\tilde{\bm{\theta}}_{\zeta}^{T}\Gamma^{-1}_{\zeta}\tilde{\bm{\theta}}_{\zeta}$
with $p\in\mathbb{R}^{+}$ being a constant, which contains $\tilde{\bm{\theta}}_{\zeta}$ rather than the full estimation error $\tilde{\bm{\theta}}$. Differentiating $V_{\zeta}$ on $t$ yields
$\dot{V}_{\zeta}=(\dot{\bm{e}}^{T}\bm{e}+\bm{e}^{T}\dot{\bm{e}})/2+(1+p){\tilde{\bm{\theta}}}_{\zeta}^{T}\Gamma^{-1}_{\zeta}\dot{\tilde{\bm{\theta}}}_{\zeta}.$
Applying (50) to the above formula, one obtains
${\dot{V}}_{\zeta}=$ $\bm{e}^{T}(-K\bm{e}+\Phi^{T}_{\zeta}{\tilde{\bm{\theta}}}_{\zeta})$
$-(1+p)({\tilde{\bm{\theta}}}_{\zeta}^{T}\Phi_{{\rm f},\zeta}\Phi_{{\rm f},\zeta}^{T}\tilde{\bm{\theta}}_{\zeta}+\kappa\tilde{\bm{\theta}}^{T}_{\zeta}Q_{\zeta}(t,t_{\rm e}){\tilde{\bm{\theta}}}_{\zeta}).$ (51)
Noting the derivation from (42) to (B), one omits the similar steps to directly give the following result:
$\dot{V}\!\leq\!-k_{\rm c}\bm{e}^{T}\bm{e}\!-\!\kappa(1\!+\!p)\tilde{\bm{\theta}}_{\zeta}^{T}Q_{\zeta}(t,t_{\rm e}){\tilde{\bm{\theta}}}_{\zeta}\!+\!\bm{e}^{T}(\Phi_{\zeta}\!-\!\Phi_{{\rm f},\zeta})\tilde{\bm{\theta}}_{\zeta}.$ (52)
As there exist some constants $T_{\rm a}$, $\sigma\in\mathbb{R}^{+}$ to satisfy the partial IE condition, one gets $\Psi_{\zeta}(t_{\rm e})\geq\sigma_{\rm c}(T_{\rm a})I\geq\sigma I$ from (18). As $H(s)$ is a stable filter with unit DC gain, there exists a constant $\sigma^{*}$ $\in$ $\mathbb{R}^{+}$ with $\sigma^{*}\leq\sigma$ such that
$Q_{\zeta}(t,t_{\rm e})=H(s)[\Psi_{\zeta}(t_{\rm e})]\geq\sigma^{*}I.$
It follows from the above result and (52) that
$\dot{V}_{\zeta}\leq-k_{\rm c}\bm{e}^{T}\bm{e}-\kappa\sigma^{*}(1+p)\tilde{\bm{\theta}}^{T}_{\zeta}\tilde{\bm{\theta}}_{\zeta}+\bm{e}^{T}(\Phi_{\zeta}-\Phi_{\rm f,\zeta})^{T}\tilde{\bm{\theta}}_{\zeta}.$
Using Lemma 2 and $\Phi$, $\Phi_{\rm f}\in L_{\infty}$, one obtains $\|\Phi-\Phi_{\rm f}\|\leq\delta$, $\forall t\geq 0$. Since the norms inactive channels are always 0, one obtains $\Phi-\Phi_{\rm f}=\Phi_{\zeta}-\Phi_{{\rm f},\zeta}$, such that
$\dot{V}_{\zeta}\leq$ $-k_{\rm c}\bm{e}^{T}\bm{e}-\kappa\sigma^{*}(1+p)\tilde{\bm{\theta}}^{T}_{\zeta}\tilde{\bm{\theta}}_{\zeta}+\delta\|\bm{e}\|\|\tilde{\bm{\theta}}_{\zeta}\|.$
Noting Young’s inequality $ab\leq a^{2}/2+b^{2}/2$ with $a=\sqrt{k_{\rm c}}\|\bm{e}\|$ and $b=\delta\|\tilde{\bm{\theta}}_{\zeta}\|/\sqrt{k_{\rm c}}$, the above inequality becomes
$\dot{V}_{\zeta}\leq$ $-k_{\rm c}\bm{e}^{T}\bm{e}/2-\kappa\sigma^{*}(1+p)\tilde{\bm{\theta}}^{T}_{\zeta}\tilde{\bm{\theta}}_{\zeta}+\delta^{2}\|\tilde{\bm{\theta}}_{\zeta}\|^{2}/(2k_{\rm c}).$
Choosing $p$ == $\delta^{2}/(2k_{\rm c}\kappa\sigma^{*})\in\mathbb{R}^{+}$, one obtains
$\dot{V}_{\zeta}\leq$ $-k_{\rm c}\bm{e}^{T}\bm{e}/2-\kappa\sigma^{*}\tilde{\bm{\theta}}^{T}_{\zeta}\tilde{\bm{\theta}}_{\zeta}$
$\leq$ $-k_{\rm c}\bm{e}^{T}\bm{e}/2-\kappa\sigma^{*}\lambda_{\min}(\Gamma_{\zeta})\tilde{\bm{\theta}}^{T}_{\zeta}\Gamma_{\zeta}^{-1}\tilde{\bm{\theta}}_{\zeta}$
$\leq$ $-k_{\rm p}V_{\zeta},t\geq T_{\rm a}$
with $k_{\rm p}$ :=:= $\min\{k_{\rm c},2\kappa\sigma^{*}\lambda_{\min}(\Gamma_{\zeta})/(1+p)\}\in\mathbb{R}^{+}$. Therefore, the equilibrium point $(\bm{e}$, $\tilde{\bm{\theta}})=\bm{0}$ of the closed-loop system (7) with (23) is partially exponentially stable on $t$ $\in$ $T_{\rm a},\infty)$, where the tracking error $\bm{e}$ and the partial estimation error $\tilde{\bm{\theta}}_{\zeta}$ exponentially converge to $\bm{0}$.
3) The proof of exponential stability under IE is similar to that under partial IE, so we omit some similar steps. Applying $V$ in (37) and the IE condition $\Psi(T_{\rm e})\geq\sigma I$ yield
$\dot{V}\leq$ $-k_{\rm c}\bm{e}^{T}\bm{e}-\kappa\sigma^{*}\lambda_{\min}(\Gamma)\tilde{\bm{\theta}}^{T}\Gamma^{-1}\tilde{\bm{\theta}}$
$\leq$ $-k_{\rm e}V,t\geq T_{\rm e}$
with $k_{\rm e}$ :=:= $\min\{k_{\rm c},2\kappa\sigma^{*}\lambda_{\min}(\Gamma)/(1+p)\}\in\mathbb{R}^{+}$ and $p$ == $\delta^{2}/$ $(2k_{\rm c}\kappa\sigma^{*})$ $\in\mathbb{R}^{+}$, where $\delta\in\mathbb{R}^{+}$ satisfies $\|\Phi-\Phi_{\rm f}\|$ $\leq$ $\delta$, $\forall t\geq 0$, and $\sigma^{*}\in\mathbb{R}^{+}$ with $\sigma^{*}\leq\sigma$ satisfies (30). Thus, the equilibrium point $(\bm{e}$, $\tilde{\bm{\theta}})=\bm{0}$ of the closed-loop system (7) with (23) is exponential stable on $t$ $\in$ $T_{\rm e},\infty)$, where the tracking error $\bm{e}$ and the estimation error $\tilde{\bm{\theta}}$ exponentially converge to $\bm{0}$. ∎
Appendix C The proof of Theorem 3
Proof.
1) The proof of Item 1 here follows the proof of Item 1 in Theorem 2. Using (25), (29), and (30), (42) becomes
$\dot{V}=$ $\bm{e}^{T}(-K\bm{e}+\Phi^{T}\tilde{\bm{\theta}})+\bm{e}^{T}\bm{d}-(1+p)\tilde{\bm{\theta}}^{T}\Phi_{{\rm f}}(\Phi_{{\rm f}}^{T}\tilde{\bm{\theta}}+{\bm{d}}_{\rm f})$
$-\kappa(1+p)\tilde{\bm{\theta}}^{T}(Q(t,t_{\rm e}){\tilde{\bm{\theta}}}+{\bm{d}}_{\rm g}).$
Following the steps from (42) to (B), one immediately gets
$\dot{V}\leq$ $-\bm{e}^{T}(K-I/4)\bm{e}-\kappa\tilde{\bm{\theta}}^{T}Q(t,t_{\rm e}){\tilde{\bm{\theta}}}+\bm{e}^{T}(\Phi-\Phi_{\rm f})\tilde{\bm{\theta}}$
$+\bm{e}^{T}\bm{d}-(1+p)\tilde{\bm{\theta}}^{T}\Phi_{{\rm f}}{\bm{d}}_{\rm f}-\kappa(1+p)\tilde{\bm{\theta}}^{T}{\bm{d}}_{\rm g}.$
Noting that the partial IE condition holds at the beginning and some moments later and using Theorem 1 and (50), the robustness of the estimation error $\tilde{\bm{\theta}}$ is ensured, thereby guaranteeing that: 1) $\tilde{\bm{\theta}}(t)\in L_{\infty}$, $\forall t\in[0,\infty)$ such that there exists a constant $c_{\theta}\in\mathbb{R}^{+}$ that satisfies $\|\tilde{\bm{\theta}}(t)\|\leq c_{\theta}$, $\forall t\in[0,\infty)$, and 2) $\bm{\epsilon}\in L_{\infty}$, $\forall t\in[0,t_{\rm f})$ such that there exists a constant $\bar{\epsilon}\in\mathbb{R}^{+}$ that satisfies $\|\bm{\epsilon}\|\leq\bar{\epsilon}$, $\forall t\in[0,t_{\rm f})$. Applying these results to the foregoing inequality and letting $k_{{\rm c}i}>1/4$ yields
$\dot{V}\leq$ $-k_{\rm c}\|\bm{e}\|^{2}+\delta c_{\theta}\|\bm{e}\|+\bar{\epsilon}(1+p)\|{\bm{d}}_{\rm f}\|$
$+\bm{e}^{T}\bm{d}-\kappa(1+p)\tilde{\bm{\theta}}^{T}{\bm{d}}_{\rm g}$
with $k_{\rm c}$ :=:= $\min_{1\leq i\leq n}\{k_{{\rm c}i}\}-1/4\in\mathbb{R}^{+}$. Noting $\|\bm{d}\|$, $\|\bm{d}_{\rm f}\|\leq\bar{d}$, and $\|{\bm{d}}_{\rm g}(t)\|$ $\leq$ ${\bar{d}}_{\rm g}$, one obtains
$\dot{V}\leq$ $-k_{\rm c}\|\bm{e}\|^{2}+\delta c_{\theta}\|\bm{e}\|+\bar{d}\|\bm{e}\|+(1+p)(\bar{\epsilon}\bar{d}+\kappa c_{\theta}{\bar{d}}_{\rm g}).$
Noting Young’s inequality $ab\leq a^{2}/2+b^{2}/2$ with $a=\sqrt{k_{\rm c}}\|\bm{e}\|$ and $b=(\delta c_{\theta}+\bar{d})/\sqrt{k_{\rm c}}$, the above inequality yields
$\dot{V}\leq-k_{\rm c}\|\bm{e}\|^{2}/2+\rho(k_{\rm c},\kappa,\bar{d})$
with $\rho$ :=:= $(\delta c_{\theta}+\bar{d})^{2}/(2k_{\rm c})+(1+p)(\bar{\epsilon}\bar{d}+\kappa c_{\theta}{\bar{d}}_{\rm g})\in\mathbb{R}^{+}$, which is valid on $\bm{x}\in\Omega_{{\rm c}_{x}}$ for $t\in[0,t_{\rm f})$. The remaining derivation is similar to the steps from (44) to the end of the proof of Item 1 in Theorem 2, so it is omitted here for saving space.
2) Consider the control problem under partial IE on $t$ $\in$ $T_{\rm a}$, $\infty)$. Using (25), (29), and (30), (B) becomes
${\dot{V}}_{\zeta}=$ $\bm{e}^{T}(-K\bm{e}+\Phi^{T}_{\zeta}{\tilde{\bm{\theta}}}_{\zeta})+\bm{e}^{T}\bm{d}-(1+p){\tilde{\bm{\theta}}}_{\zeta}^{T}\Phi_{{\rm f},\zeta}(\Phi_{{\rm f},\zeta}^{T}\tilde{\bm{\theta}}_{\zeta}+{\bm{d}}_{\rm f})$
$-\kappa(1+p)\tilde{\bm{\theta}}^{T}_{\zeta}(Q_{\zeta}(t,t_{\rm e}){\tilde{\bm{\theta}}}_{\zeta}+{\bm{d}}_{\rm g}).$
As the partial IE condition $\Psi_{\zeta}(t_{\rm e})\geq\sigma_{\rm c}(T_{\rm a})I\geq\sigma I$ is satisfied with some constants $T_{\rm a}$, $\sigma\in\mathbb{R}^{+}$, noting the proof of Item 2 in Theorem 2, the above result leads to
$\dot{V}_{\zeta}\leq$ $-k_{\rm c}\bm{e}^{T}\bm{e}-\kappa(1+p)\sigma^{*}\tilde{\bm{\theta}}^{T}_{\zeta}\tilde{\bm{\theta}}_{\zeta}+\delta\|\bm{e}\|\|\tilde{\bm{\theta}}_{\zeta}\|$
$+\bm{e}^{T}\bm{d}-(1+p){\tilde{\bm{\theta}}}_{\zeta}^{T}\Phi_{{\rm f},\zeta}{\bm{d}}_{\rm f}-\kappa(1+p){\tilde{\bm{\theta}}}_{\zeta}^{T}{\bm{d}}_{\rm g}.$
Noting $\|\bm{d}\|$, $\|\bm{d}_{\rm f}\|\leq\bar{d}$, and $\|{\bm{d}}_{\rm g}(t)\|$ $\leq$ ${\bar{d}}_{\rm g}$, one obtains
$\dot{V}_{\zeta}\leq$ $-k_{\rm c}\|\bm{e}\|^{2}-\kappa(1+p)\sigma^{*}\|\tilde{\bm{\theta}}_{\zeta}\|^{2}+\delta\|\bm{e}\|\|\tilde{\bm{\theta}}_{\zeta}\|$
$+\bar{d}\|\bm{e}\|+(1+p)(\bar{\epsilon}\bar{d}+\kappa{\bar{d}}_{\rm g}\|\tilde{\bm{\theta}}_{\zeta}\|).$
Following the proof of Item 2 in Theorem 2, one obtains
$\dot{V}_{\zeta}\leq$ $-k_{\rm c}\|\bm{e}\|^{2}/2+\bar{d}\|\bm{e}\|-\kappa\sigma^{*}\|\tilde{\bm{\theta}}_{\zeta}\|^{2}$
$+(1+p)\kappa{\bar{d}}_{\rm g}\|\tilde{\bm{\theta}}_{\zeta}\|+(1+p)\bar{\epsilon}\bar{d}$
with $p$ == $\delta^{2}/(2k_{\rm c}\kappa\sigma^{*})\in\mathbb{R}^{+}$, which can be rewritten into
$\dot{V}_{\zeta}\leq$ $-k_{\rm c}(1-\mu)\|\bm{e}\|^{2}/2-k_{\rm c}\mu(\|\bm{e}\|^{2}-2\bar{d}/(k_{\rm c}\mu)\|\bm{e}\|)/2$
$-\kappa\sigma^{*}\mu(\|\tilde{\bm{\theta}}_{\zeta}\|^{2}-(1+p){\bar{d}}_{\rm g}/(\sigma^{*}\mu)\|\tilde{\bm{\theta}}_{\zeta}\|)$
$-\kappa\sigma^{*}(1-\mu)\|\tilde{\bm{\theta}}_{\zeta}\|^{2}+(1+p)\bar{\epsilon}\bar{d}$
with $\mu\in(0,1)$. Applying Young’s inequality $2ab-a^{2}\leq b^{2}$ with $a=\|\bm{e}\|$ or $\|\tilde{\bm{\theta}}_{\zeta}\|$ and $b=\bar{d}/(k_{\rm c}\mu)$ or $(1+p){\bar{d}}_{\rm g}/(2\sigma^{*}\mu)$ to the above expression, one gets
$\dot{V}_{\zeta}\leq$ $-k_{\rm c}(1-\mu)\|\bm{e}\|^{2}/2-\kappa\sigma^{*}(1-\mu)\|\tilde{\bm{\theta}}_{\zeta}\|^{2}+\rho_{2}(k_{\rm c},\kappa,\bar{d})$
with $\rho_{2}(k_{\rm c},\kappa,\bar{d})$ :=:= $(1+p)\bar{\epsilon}\bar{d}+\bar{d}^{2}/(2k_{\rm c}\mu)+\kappa((1+p){\bar{d}}_{\rm g})^{2}$ $(4\sigma^{*}\mu) \in$ $\mathbb{R}^{+}$. It follows from the above result that
$\dot{V}_{\zeta}\leq$ $-k_{\rm c}(1-\mu)\bm{e}^{T}\bm{e}/2-\kappa\sigma^{*}\lambda_{\min}(\Gamma_{\zeta})(1-\mu)\tilde{\bm{\theta}}_{\zeta}^{T}\Gamma^{-1}_{\zeta}\tilde{\bm{\theta}}_{\zeta}$
$+\rho_{2}(k_{\rm c},\kappa,\bar{d}).$
Consequently, one immediately gets
$\dot{V}_{\zeta}\leq$ $-k_{\rm r}V_{\zeta}+\rho_{2}(k_{\rm c},\kappa,\bar{d}),\forall t\geq T_{\rm a}$
with $k_{\rm r}$ :=:= $\min\{k_{\rm c}(1-\mu),2\kappa\sigma^{*}\lambda_{\min}(\Gamma_{\zeta})(1-\mu)/(1+p)\}\in\mathbb{R}^{+}$. Applying [88, Lemma A.3.2] to the above inequality yields
$V_{\zeta}(t)\leq(V_{\zeta}(0)-\rho_{2}(k_{\rm c},\kappa,\bar{d}))e^{-k_{\rm r}t}+\rho_{2}(k_{\rm c},\kappa,\bar{d}),\forall t\geq T_{\rm a}$
which impiles that the equilibrium point $(\bm{e}$, $\tilde{\bm{\theta}})=\bm{0}$ of the closed-loop system (25) with (23) has partial practical exponential stability on $t$ $\in$ $T_{\rm a},\infty)$, where the tracking error $\bm{e}$ and the partial estimation error $\tilde{\bm{\theta}}_{\zeta}$ exponentially converge to a small neighborhood of $\bm{0}$ dominated by $k_{{\rm c}i}$, $\kappa$, and $\bar{d}$.
3) The proof of robustness under IE is similar to that under partial IE, so we omit some similar steps. Applying $V$ in (37) and the IE condition $\Psi(T_{\rm e})\geq\sigma I$, one obtains
$\dot{V}\leq$ $-k_{\rm c}(1-\mu)\bm{e}^{T}\bm{e}/2-\kappa\sigma^{*}\lambda_{\min}(\Gamma)(1-\mu)\tilde{\bm{\theta}}^{T}\Gamma^{-1}\tilde{\bm{\theta}}+\rho_{2}$
$\leq$ $-k_{\rm r}V+\rho_{2}(k_{\rm c},\kappa,\bar{d}),\forall t\geq T_{\rm e}$
with $k_{\rm r}$ :=:= $\min\{k_{\rm c}(1-\mu),2\kappa\sigma^{*}\lambda_{\min}(\Gamma)(1-\mu)/(1+p)\}\in\mathbb{R}^{+}$, $\rho_{2}$ :=:= $(1+p)\bar{\epsilon}\bar{d}+\bar{d}^{2}/(2k_{\rm c}\mu)+\kappa((1+p){\bar{d}}_{\rm g})^{2}/(4\sigma^{*}\mu)\in\mathbb{R}^{+}$, and $p$ == $\delta^{2}/(2k_{\rm c}\kappa\sigma^{*})\in\mathbb{R}^{+}$, where $\delta\in\mathbb{R}^{+}$ satisfies $\|\Phi-\Phi_{\rm f}\|$ $\leq$ $\delta$, $\forall t\geq 0$, and $\sigma^{*}\in\mathbb{R}^{+}$ with $\sigma^{*}\leq\sigma$ satisfies (30). Thus, the equilibrium point $(\bm{e}$, $\tilde{\bm{\theta}})=\bm{0}$ of the closed-loop system (25) with (23) is practical exponential stable on $t$ $\in$ $T_{\rm e},\infty)$, where the tracking error $\bm{e}$ and the estimation error $\tilde{\bm{\theta}}$ exponentially converge to a small neighborhood of $\bm{0}$ dominated by $k_{{\rm c}i}$, $\kappa$, and $\bar{d}$. ∎